### v1

# Learning Loss Landscapes in Preference Optimization

Abstract
We present an empirical study investigating how specific properties of preference datasets, such as mixed-quality or noisy data, affect the performance of Preference Optimization (PO) algorithms. Our experiments, conducted in MuJoCo environments, reveal several scenarios where state-of-the-art PO methods experience significant drops in performance. To address this issue, we introduce a novel PO framework based on mirror descent, which can recover existing methods like Direct Preference Optimization (DPO) and Odds-Ratio Preference Optimization (ORPO) for specific choices of the mirror map. Within this framework, we employ evolutionary strategies to discover new loss functions capable of handling the identified problematic scenarios. These new loss functions lead to significant performance improvements over DPO and ORPO across several tasks. Additionally, we demonstrate the generalization capability of our approach by applying the discovered loss functions to fine-tuning large language models using mixed-quality data, where they outperform ORPO.
1 Introduction
Learning from human feedback is a paradigm that enables the alignment of complex agents to human preferences, and has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023). In particular, fine-tuning pretrained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
Most LLM alignment pipelines begin with a supervised fine-tuning (SFT) step, which involves supervised next-token prediction on a dataset of high-quality responses and leads to a reference policy. The reference policy is further optimized using the human preference data, typically through either Reinforcement Learning from Human Feedback (RLHF)  (Christiano et al., 2017) or Direct Preference Optimization (DPO)  (Rafailov et al., 2024), or one of their several variants. RLHF consists of learning a reward model consistent with human preferences and then using Reinforcement Learning (RL) techniques such as REINFORCE (Sutton et al., 1999) and Proximal Policy Optimisation (PPO)  (Schulman et al., 2017) to maximize the total expected reward. In contrast, DPO and its variations, e.g. odds ratio preference optimization (ORPO) (Hong et al., 2024), bypass explicit reward models entirely and optimize directly on preference data, implicitly learning the reward.
While RL-based methods offer stronger theoretical guarantees and often lead to higher performance (Song et al., 2024; Xu et al., 2024), offline approaches such as DPO have gained traction due to their simplicity and the ability to leverage preexisting high-quality datasets. In contrast to PPO, where data collection and labeling are performed iteratively after each update, DPO and its modifications allow for more efficient training, avoiding the high computational costs, need of additional sample labelling and complexity of RL methods. Specifically, PPO requires careful parameter tuning (Yuan et al., 2023) and involves simultaneous training of multiple models (the reward model, language model, and critic), which can be prohibitive in most hardware setups.
The performance of offline algorithms such as DPO, particularly on noisy or low-quality datasets, has been a subject of debate (Chowdhury et al., 2024), with limited empirical results available. In this work, we provide a comprehensive analysis of PO algorithms, examining their behavior on automatically generated preference datasets. We perform this analysis in MuJoCo environments (Todorov et al., 2012), where the underlying reward structure is well defined and offers a clear performance metric to compare agents. In particular, we focus on ORPO, as we have found that it largely outperforms DPO in all settings we have considered. Our findings indicate that ORPO exhibits distinct failure modes when applied to specific low-quality or noisy datasets. These failure modes are present in practical LLM applications, raising concerns about using mixed-quality datasets for PO.
To find algorithms that are capable of dealing with these more difficult settings, we introduce a novel framework for PO based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO and allows to search for new Preference Optimization (PO) algorithms. We perform this search over the space of mirror maps using evolutionary strategies (ES), as they have been shown to be effective for discovering algorithms (Lu et al., 2022; Jackson et al., 2024). We outline our contributions in detail below.
  1. 1.
    We perform a systematic analysis of ORPO on preference datasets with varying levels of data quality, noise levels, initial policy, and judge temperature.
  2. 2.
    We introduce a novel family of offline PO algorithms using mirror descent, which can be easily parameterized and which we explore via ES.
  3. 3.
    For each failure mode we identify, we find and describe an algorithm within our framework that outperforms ORPO. Additionally, we show that allowing losses to vary based on the percentage of training progress drastically boosts performance in some settings.
  4. 4.
    We demonstrate that PO algorithms discovered on MuJoCo can generalize to LLM tasks. In particular, for a highlighted failure setting, we show that our discovered algorithm improves performance w.r.t. the ORPO baseline. This finding showcases the efficacy of analyzing PO algorithms in simpler, less computationally expensive environments, as it is possible to obtain insights relevant for LLM applications.
2 Preliminaries
Let $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,T,\mu)$ denote an episodic Markov Decision Process, where $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action spaces, $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ when taking action $a$, $r(s,a)\in[0,1]$ is the reward function, $T$ is the maximum episode length, and $\mu$ is a starting state distribution. A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$. Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$. Lastly, let $\pi(\cdot|\tau)$ be the product distribution of $\pi(\cdot|s_{0}),\dots,\pi(\cdot|s_{N-1})$.
Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$. Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$. We assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
$\mathbb{P}(\tau_{w}\succ\tau_{l})=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$ (1)
where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter. Our objective is to exploit the dataset $\mathcal{D}$ to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau):=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\sum_{t=0}^{T-1}r(s_{t},a_{t}).$ (2)
We consider an offline setting, where we do not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
2.1 Preference optimization
There are several methods in the literature to optimize the objective in (2) using a preference dataset $\mathcal{D}$. We provide here a short review of the main pipelines (Ouyang et al., 2022). Two common independent preliminary steps are reward modelling (RM) and supervised fine tuning (SFT). RM aims to obtain an estimate of the true reward function, and is usually framed as a maximum likelihood estimation problem for a Bradley-Terry preference model, i.e. find
$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}\log\sigma(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})/\eta),$ (3)
for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$ and for $\eta\geq 0$. SFT is an initial alignment phase, where the starting policy $\pi_{0}$ is trained to imitate high-quality demonstration data. In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$, typically doing one epoch over the dataset. We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure. The new objective we want to optimize is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$ (4)
where the $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution. The expressions in (3) and (4) can be optimized sequentially, first obtaining a reward estimate in (3) and then optimizing (4) with PPO using the reward estimate in place of the true reward. Alternatively, the optimization problems in (3) and (4) can be solved implicitly using DPO or ORPO. We proceed to discuss each of these methods.
PPO
PPO is recognized as one of the preferred methods for optimizing (4) when the necessary computing resources are available, as demonstrated by its success in training state of the art models like GPT-4 (Achiam et al., 2023) and Claude (Antropic, 2023). However, it presents a complex pipeline where one needs to effectively train a reward model, perform SFT and then optimize (4), where each phase has a different set of hyper-parameters. Additionally, storing the reward model, the reference agent and the current agent in memory is impractical in most setups and often requires sacrificing other aspects, such as batch-size. Besides it computational costs, PPO is known to be prone to reward overoptimization (Coste et al., 2024).
DPO
Direct Preference Optimization (DPO) bypasses the need for an explicit reward model by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$ (5)
which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3). Refer to Appendix A for details. Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs as an alternative to PPO (Yuan et al., 2024; Jiang et al., 2024).
A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses, which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences. In contrast, PPO can leverage the generalization capabilities of the (learned) reward model to generate responses beyond the preference dataset distribution, while the $\mathtt{KL}$-divergence penalty can provide additional regularization. To mitigate the risks described above, DPO is usually only applied for a few epochs.
ORPO
ORPO is a more recent algorithm that aims to further simplify the training pipeline and, concurrently, to address the distribution shift issue present in DPO. ORPO merges the SFT and DPO steps into one, which optimizes the unified objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\right]$ (6)
where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$. ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, which substitutes the role of the SFT step and of the reference model in preventing the optimized policy from moving too far away from the dataset distribution. Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.
2.2 Mirror Maps
We review the concept of mirror map, which will be needed when describing our methodology. For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function [^1] A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$. function that satisfies $\nabla h(\mathcal{X})=\mathbb{R}^{|\mathcal{A}|}$. Essentially, a mirror map is a function whose gradient allows bijective mapping between the primal space $\mathcal{X}$ and the dual space $\mathbb{R}^{|\mathcal{A}|}$. The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
Definition 2.1 ($\omega$-potential mirror map Krichene et al. (2015)).
For $u\in(-\infty,+\infty]$, $\omega\leq 0$, an $\omega$-potential is defined as an increasing $C^{1}$-diffeomorphism $\phi:(-\infty,u)\rightarrow(\omega,+\infty)$ such that
[…]
For any $\omega$-potential $\phi$, we define the associated mirror map $h_{\phi}$ as
$h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix B). Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one. Additionally, the same $\omega$-potential $\phi$ can be used to generate mirror maps for different action spaces, allowing the insights obtained for one action space to easily generalize to others. An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s)):=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$. When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix B). When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
3 Methodology
We develop a new framework for preference optimization based on mirror maps, which generalizes DPO and ORPO. We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a more general Bregman divergence, that is, we aim to solve the problem
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$ (7)
where $D_{h}$ is the Bregman divergence induced by a mirror map $h$. This new objective allows us to enforce different types of regularization, which, as we show later in the paper, can be tailored to account for specific properties of the preference dataset. Following the same intuition used to obtain the DPO objective, we have the following result.
Theorem 3.1.
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
$r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (8)
where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
We provide a proof for Theorem 3.1 in Appendix A. By plugging (8) in the maximum likelihood problem in (3), we obtain the objective:
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}\log\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$ (9)
where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$. When $\phi=e^{x}$, the expression in (7) recovers the DPO objective in (5). Additionally, we can modify the expression in (7) to include an SFT term in order to avoid the SFT step, in a manner similar to ORPO. Let $\psi$ be an $\omega$-potential and let $\pi_{\mathrm{ref}}$ be the uniform distribution, then the new objective we consider is
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\psi(\pi(\tau_{w}))+\lambda\log\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\right],$ (10)
where the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ do not appear as they have cancelled out. We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization. When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x)-\log(1-x)$, (10) recovers the ORPO objective in (6).
The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification. In the following, we will show that taking into account the properties of the preference dataset when choosing $\psi$ and $\phi^{-1}$ can lead to a better performance of the trained policy.
Following the intuition developed by Jackson et al. (2024), we also allow our objective function, specifically $\psi$ and $\phi^{-1}$, to account for training progress. This is equivalent to changing the Bregman divergence penalty in (7) during training and, as we shall discuss in the following section, helps in dealing with datasets with mixed-quality data.
3.1 Learning mirror maps
To search the space of PO algorithms we have defined, we employ a neural network parametrization for both $\psi$ and $\phi^{-1}$, which we optimize using evolutionary strategies.
Similarly to Alfano et al. (2024), we parameterize both $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),\;e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$. The non-negative kernels and the increasing activation functions guarantee the monotonicity of $\psi$ and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions. To ensure that we are able to recover the ORPO objective, we add $a\log(x)$ and $b\log(x)-b\log(1-x)$ to the final outputs of $\psi$ and $\phi^{-1}$, respectively, where $a,b\geq 0$. In case we want to take into account training progress, we give a second input to the neural network, i.e. $x\cdot n/N$, where $n$ is the current epoch and $N$ is the total number of epochs through the dataset. To ensure that monotonicity is preserved, we lower bound the weights associated to the second input with the negative of the respective weights associated to the first input.
To search for the best $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy (Salimans et al., 2017). Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient.
4 Experiments
We carry out our experiments on continuous reinforcement learning tasks in MuJoCo and on LLM fine-tuning. In particular, due to computational efficiency constraints, we discover mirror maps in MuJoCo and report their performance when evaluated for an LLM finetuning task.
4.1 Environments
We consider three different preference optimization settings which we describe below. Each setting includes a starting policy, a set of preference datasets with different properties and an evaluation method for the trained policy. We apply ES to the first two settings and test the discovered objectives on the last one.
Hopper
Our objective in Hopper is to learn a policy from scratch, i.e. with a randomly initialised policy, using a preference dataset. On Hopper, we train four reference agents of different skill levels, specifically, of respective expected cumulative reward (or value) of 900, 1200, 1800, and 2100 (the expert agent). Each dataset consists of 5120 rows, each with two trajectories of length 1000 starting from the same state. We generate a dataset by comparing trajectories by the expert agent (average return 2100), with other trajectories by one of our reference agents. A Bradley-Terry judge ranks each pair of trajectories, based on their true reward. The trained policy is then evaluated on the standard Hopper environment. To test the efficacy and robustness of our loss functions, we analyse a few variations of these standard datasets that represent common issues of real world data.
  • Base dataset: we compare a trajectory from the expert agent with one of a reference agent.
  • Noisy dataset: we flip a varying portion of the rankings in the base dataset.
  • Shuffled dataset: the comparisons are not always between the expert agent and a second agent. Namely, $25\%$ of the comparisons are between two trajectories from the expert agent, $50\%$ of the comparisons are between a trajectory of the expert agent and one of a reference agent, and $25\%$ of the comparisons are between two trajectories of a reference agent.
  • Poor judge dataset: the judge is more likely to flip labels when the two trajectories have closer reward values. This is implemented as an increase in the temperature of the Bradley-Terry judge.
Ant
To further investigate a scenario similar to fine-tuning a large language model (LLM), we examine a setting in which a pre-trained agent is required to modify its behavior to achieve its goal with a stylistic constraint. Specifically, in the Ant environment, we take an agent that has been pre-trained on the standard Ant goal of moving forward, and enforce the objective of avoiding the use of one of its legs. This is accomplished by introducing the Three-legged-ant (TLA) environment, a modified version of Ant where utilizing the fourth leg results in significant penalties. We train one agent in the original Ant environment, achieving a reward of 6000, and another in the TLA environment, which achieves a reward of 3900. For comparison, the agent trained in the original Ant environment achieves a reward of 1700 when evaluated in the TLA environment. The dataset generation follows the same protocol as described for the Hopper environment, where trajectories are collected from both the Ant and TLA agents. However, in this setting the number of rows for each dataset is 1280, to account for the fact that we do not want the agent to learn from scratch, but to adjust its policy to a new instruction. The trained policy is then evaluated on the TLA environment.
LLM tuning
Finally, we evaluate one of our discovered objective functions on a real-world LLM fine-tuning task. To simulate a scenario involving mixed data quality, we fine-tune the gemma-7b [^2] https://huggingface.co/google/gemma-7b model on a modified version of the dpo-mix-7k [^3] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset, where half of the responses, selected at random, are replaced with responses generated by gemma-2b [^4] https://huggingface.co/google/gemma-2b, a model that typically produces lower-quality responses compared to those originally present in DPO-mix. This approach replicates the shuffled dataset described for the MuJoCo experiments, aiming to simulate the challenges of training with datasets of varying quality, a common issue in fine-tuning tasks where it is difficult and costly to ensure uniformly high-quality data.
We implement our experiments in MuJoCo using the brax library (Freeman et al., 2021) for the environments and the evosax library (Lange, 2022) for the evolution. We use the same hyper-parameters for all MuJoCo datasets, which have been tuned for ORPO using Weights and Biases [^5] https://wandb.ai/site. We report them in Appendix C, along the hyper-parameters for OpenES. For the LLM tuning, we modify the Alignment Handbook library (Tunstall et al., ), and use the default hyper-parameters.
4.2 Results: Mujoco
Table 1: Results for Hopper. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset. When ranking two trajectories of value 2100 and 1800, a judge with low, medium, and high temperature provides the correct ranking 95%, 85% and 75% of the time, respectively.
Value agent 1 Value agent 2 Noise Shuffled Judge temp. ORPO Discovered
2100 900 0 No Low 2003$\pm$20 2012$\pm$11
2100 1200 0 No Low 2055$\pm$12 2055$\pm$10
2100 1800 0 No Low 2043$\pm$ 15 2073$\pm$12
2100 2100 0 No Low 2070$\pm$12 2098$\pm$11
2100 900 0.1 No Low 1519$\pm$45 1724$\pm$43
2100 1800 0.1 No Low 1936$\pm$25 1945$\pm$25
2100 2100 0.1 No Low 2112$\pm$9 2094$\pm$13
2100 900 0.2 No Low 662$\pm$49 1127$\pm$45
2100 900 0.3 No Low 623$\pm$29 698$\pm$26
2100 1800 0 No Medium 1884$\pm$22 1928$\pm$23
2100 1800 0 No High 1858$\pm$32 1895$\pm$14
2100 900 0 Yes Low 975$\pm$51 1099$\pm$35
Table 2: Results for TLA. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset.
Value agent 1 Value agent 2 Noise Shuffled ORPO Discovered
3900 1700 0 No 3277$\pm$49 3473$\pm$63
3900 1700 0 Yes 2425$\pm$58 3485$\pm$163
3900 1700 0.1 No 2675$\pm$62 3837$\pm$52
Table 3: Results for temporally-aware objectives. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the temporally-aware objective discovered for that dataset.
Value agent 1 Value agent 2 Env Noise Shuffled ORPO Discovered
2100 900 Hopper 0 Yes 975$\pm$51 2032$\pm$13
3900 1700 TLA 0 No 3277$\pm$49 3903$\pm$42
3900 1700 TLA 0 Yes 2425$\pm$58 3251$\pm$83
We provide the results of our experiments for MuJoCo in Tables 12, and 3, which report the performance of ORPO and our discovered objectives for Hopper, TLA, and the temporally aware case, respectively. To provide information on the behavior of these objectives as well as guidelines on how to design loss landscapes for the settings we have considered, we compare the landscapes of ORPO and of the discovered objectives in Figures 123, and 4. In these figures, we report, in absolute value, the gradients of the objective with respect to $\log(\pi(\tau_{w}))$ and $\log(\pi(\tau_{l}))$, as well as a sampled training trajectory. We discuss the results below, by dataset type.
Base Dataset
For the base dataset, we observe different behaviors for the Hopper and TLA environments. For the Hopper dataset, we see that ORPO achieves the performance of the best reference agent in all cases and that it is not significantly outperformed by the learned objective. On the other hand, ORPO does not reach the performance of the best reference agent in the TLA dataset and is surpassed by the discovered objectives. Remarkably, the temporally aware objective achieves the performance of the best reference agent, presenting a $19\%$ improvement over ORPO. As shown by Figures 1 and  4, both the static and the temporally aware discovered objectives present a smaller gradient than ORPO, which means that they cause smaller updates. Additionally, for ORPO we have that if $\log(\pi(\tau_{w}))<\log(\pi(\tau_{l}))$ then the gradients increases suddenly by half, while for the temporally aware discovered objective we can observe that the gradient increases more slowly. Essentially, this means that ORPO causes a large update when the agent mistakenly assigns more probability to the rejected trajectory, while the temporally aware discovered objective causes a large update only if the agent assigns significantly more probability to the rejected trajectory.
Figure 1: The plots represent the gradients, in absolute value, of ORPO (left) and the objective discovered on the base TLA dataset (right). The dots represent a sampled training trajectory.
Noisy Dataset
As expected, the performance of ORPO declines when the dataset contains randomly flipped labels, with the degradation becoming more pronounced as the percentage of flipped labels increases. While the performance degrades for our discovered objectives in Hopper as well, it happens more slowly and we show significant improvements over ORPO for all noise levels. In TLA, the discovered objective is capable of almost achieving the best reference agent value despite the noise. Figure 2 shows that the objective discovered in Hopper assigns a roughly constant value to the gradient w.r.t. the chosen logits, while the gradient w.r.t. the rejected logits increases as $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ decreases. Consequently, when the agent encounters a flipped label where $\log(\pi(\tau_{w}))\ll\log(\pi(\tau_{l}))$, the agent does not take larger updates to increase the chosen logits, but only larger updates to decrease the rejected logit, thus mitigating the noise effects. On the other hand, the objective discovered in TLA manages to mitigate the influence of noise by reducing the gradient value and by taking larger updates only when the chosen logits are much larger than the rejected logits.
Figure 2: The plots represent the gradients, in absolute value, of the objectives discovered on the noisy Hopper dataset (left) and on the noisy TLA dataset (right). The level of noise is 0.1 for both plots. The dots represent a sampled training trajectory.
Shuffled Dataset
While effectively using the same data as the base dataset, the performance of ORPO significantly decreases in both the Hopper and TLA shuffled dataset, where low quality trajectories are selected as the preferred choice when compared to other low quality trajectories. In contrast, the discovered objectives achieve a value of the final policy close to the best reference agent. In particular, the value reached by the temporally aware objective for Hopper is more than double the value reached by ORPO. All discovered objectives present smaller gradients than ORPO, and therefore induce smaller updates, which prevents taking large updates when the chosen logits belong to a low-quality trajectory. As to the temporally aware objective discovered on HopperFigure 4 shows that it induces very small updates for the chosen logits when they are low, and therefore probably belong to a low-quality trajectory. Lastly, we note a decrease in the absolute value of the gradient at the top and right borders as training progresses, which induce smaller updates.
Figure 3: The plots represent the gradients, in absolute value, of the objectives discovered on the shuffled Hopper dataset (left) and on the shuffled TLA dataset (right). The dots represent a sampled training trajectory.
Poor judge
As expected, we observe a decrease in performance for ORPO when we increase the temperature of the judge. However, since we compared trajectories from two high-quality reference agents, the drop in performance is small. This is also reflected in the discovered objectives, which do not manage to significantly outperform ORPO.
Figure 4: The plots represent the gradients, in absolute value, of the temporally aware objectives discovered on the shuffled Hopper dataset (left) and on the base TLA dataset (right), for three different values of the training progress. For each row, the dots represent the respective third of a sampled training trajectory.
4.3 Results: Transfer
We show that PO objectives learnt with our methodology can transfer to other domains. In particular, the static objective discovered on the shuffled TLA dataset tranfers to both the shuffled Hopper dataset, where it obtains a final policy value of 1735$\pm$39, and to the LLM-tuning task with shuffled data defined above. We train the base LLM on the modified DPO-mix dataset with both ORPO and the discovered objective. We observe that the first methodology achieves 57% accuracy on the test set while the latter achieves 62%. Additionally, we compare the two trained models with AlpacaEval (Li et al., 2023), and obtain a 53% winrate for the model trained with the discovered objective.
5 Related Work
Automatic Discovery of Preference Optimization Loss Functions
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by Lu et al. (2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by Tang et al. (2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
Generalisations of DPO
A generalization of DPO alternative to ours is $f$-DPO, developed by Wang et al. (2023), which consists in replacing the $\mathtt{KL}$-divergence in (2) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. Huang et al. (2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
6 Conclusion
We have introduced a novel framework for Policy Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolution strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we demonstrated that our discovered objective functions consistently match or exceed the performance of existing methods, particularly in noisy and mixed-quality datasets where the ORPO baseline struggles. The introduction of temporally aware objective functions further improved performance, allowing the optimization process to vary and adapt during training. Analysing the landscape of the discovered objectives, we give an intuition that justifies improved performance as well as guidance in the design of new PO algorithms. Our results also indicate that the discovered objectives generalize beyond simple reinforcement learning environments, showing promising performance when transferred to LLMs, thereby confirming the broader applicability of our approach.
Appendix A Proof of Theorem 3.1
We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follow by taking $\phi=e^{x}$
Theorem A.1 (Theorem 3.1).
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
$r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (11)
for all trajectories $\tau$, where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
Proof.
We use the KKT conditions to solve (7), i.e.
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right]$
We use the stationarity condition to obtain the equation
$\nabla_{\pi(\tau)}$ $\Bigg{[}\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))$
$\quad-\lambda\left(\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\pi(\tau^{\prime})-1\right)+\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg{]}$
$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
for all initial states $s_{0}\in\mathcal{S}$ and for all trajectories $\tau$ starting from $s_{0}$. Rearranging, we obtain that
$\pi(\tau)=\phi\big{(}r(\tau)+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)\big{)}.$
Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 00-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$. Invoking the complementary slackness condition, whereby $\alpha(\tau)\pi(\tau)=0$ for all trajectories $\tau$, we have that $\alpha(\tau)=0$ for all trajectories $\tau$. Therefore, we have that
$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
The theorem statement is obtained by rearranging the last equation and denoting $c(s_{0})=\lambda$
Appendix B Further discussion of $\omega$-potentials
We show here two examples of Bregman divergence induced by an $\omega$-potential mirror map, that is when $\phi(x)=e^{x-1}$ and when $\phi(x)=x$. If $\phi(x)=e^{x-1}$, the associated mirror map is defined as
[…]
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)\log(\pi(a\mid s))-\pi(a\mid s)+\pi(a\mid s)$
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)\log(\pi(a\mid s)),$
which is the negative entropy. Plugging this expression in the definition of Bregman divergence we obtain
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
$=\sum_{a\in\mathcal{A}}x_{a}\log(x_{a})-y_{a}\log(y_{a})-(\log(y_{a})-y_{a})(x_{a}-y_{a})$
$=\sum_{a\in\mathcal{A}}x_{a}\log(x_{a}/y_{a}),$
which is the definition of the $\mathtt{KL}$-divergence. If $\phi(x)=2x$, the associated mirror map is defined as
[…]
$=\sum_{a\in\mathcal{A}}\pi(a\mid s)^{2},$
which is the $\ell_{2}$-norm. Plugging this expression in the definition of Bregman divergence we obtain
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
$=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})$
$=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
which is the definition of the Euclidean distance.
Appendix C Hyper-parameters
We give the hyper-parameters we use for training in Tables 4 and 5.
Parameter Value
Number of epochs 12
Minibatch size 2
Learning rate 1e-3
Max gradient norm 1.3
$\lambda$ 0.5
Table 4: Hyper-parameter settings for PO.
Parameter Hopper TLA
Population Size 256 256
Number of generations 128 256
Sigma init 0.03 0.03
Sigma Decay 0.999 0.999
Learning rate 0.02 0.02
Table 5: Hyper-parameter settings of OpenAI-ES.

### v2

@@
-Learning Loss Landscapes in Preference Optimization
+Meta-Learning Objectives for Preference Optimization
 Abstract
-We present an empirical study investigating how specific properties of preference datasets, such as mixed-quality or noisy data, affect the performance of Preference Optimization (PO) algorithms.
-Our experiments, conducted in MuJoCo environments, reveal several scenarios where state-of-the-art PO methods experience significant drops in performance.
-To address this issue, we introduce a novel PO framework based on mirror descent, which can recover existing methods like Direct Preference Optimization (DPO) and Odds-Ratio Preference Optimization (ORPO) for specific choices of the mirror map.
-Within this framework, we employ evolutionary strategies to discover new loss functions capable of handling the identified problematic scenarios.
-These new loss functions lead to significant performance improvements over DPO and ORPO across several tasks.
-Additionally, we demonstrate the generalization capability of our approach by applying the discovered loss functions to fine-tuning large language models using mixed-quality data, where they outperform ORPO.
+Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters.
+In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on much simpler benchmarks.
+We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark.
+We then propose a novel family of PO algorithms based on mirror descent, which we call Mirror Preference Optimization (MPO).
+Through evolutionary strategies, we search this class to discover algorithms specialized to specific properties of preference datasets, such as mixed-quality or noisy data.
+We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings.
+Finally, based on the insights gained from our MuJoCo experiments, we design a novel PO algorithm that significantly outperforms existing baselines in an LLM alignment task.
 1 Introduction
-Learning from human feedback is a paradigm that enables the alignment of complex agents to human preferences, and has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023).
-In particular, fine-tuning pretrained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
-Most LLM alignment pipelines begin with a supervised fine-tuning (SFT) step, which involves supervised next-token prediction on a dataset of high-quality responses and leads to a reference policy.
-The reference policy is further optimized using the human preference data, typically through either Reinforcement Learning from Human Feedback (RLHF)  (Christiano et al., 2017) or Direct Preference Optimization (DPO)  (Rafailov et al., 2024), or one of their several variants.
-RLHF consists of learning a reward model consistent with human preferences and then using Reinforcement Learning (RL) techniques such as REINFORCE (Sutton et al., 1999) and Proximal Policy Optimisation (PPO)  (Schulman et al., 2017) to maximize the total expected reward.
-In contrast, DPO and its variations, e.g.
-odds ratio preference optimization (ORPO) (Hong et al., 2024), bypass explicit reward models entirely and optimize directly on preference data, implicitly learning the reward.
-While RL-based methods offer stronger theoretical guarantees and often lead to higher performance (Song et al., 2024; Xu et al., 2024), offline approaches such as DPO have gained traction due to their simplicity and the ability to leverage preexisting high-quality datasets.
-In contrast to PPO, where data collection and labeling are performed iteratively after each update, DPO and its modifications allow for more efficient training, avoiding the high computational costs, need of additional sample labelling and complexity of RL methods.
-Specifically, PPO requires careful parameter tuning (Yuan et al., 2023) and involves simultaneous training of multiple models (the reward model, language model, and critic), which can be prohibitive in most hardware setups.
-The performance of offline algorithms such as DPO, particularly on noisy or low-quality datasets, has been a subject of debate (Chowdhury et al., 2024), with limited empirical results available.
+Learning from human preferences (Christiano et al., 2017) is a paradigm which enables the alignment of machine learning systems to relative human preferences, without requiring access to absolute rewards.
+While the framework was developed for robotic and games applications with experiments on MuJoCo simulations and Atari (Akrour et al., 2012; Biyik & Sadigh, 2018; Ibarz et al., 2018), this paradigm has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023).
+In particular, fine-tuning pre-trained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
+Within this framework, Reinforcement Learning from Human Feedback (RLHF) is one of the most popular methods.
+It consists in learning a reward function using a preference dataset and then optimizing the estimated reward using Reinforcement Learning methods such as Proximal Policy Optimization (PPO)  (Schulman et al., 2017).
+However, the training pipeline of RLHF is quite complex, which is why implicit approaches such as Direct Preference Optimisation (DPO)  (Schulman et al., 2017) have gained traction thanks to their simplicity.
+These methods do not learn a reward model but estimate it implicitly using the policy of the agent.
+Many follow ups to DPO have been proposed  (Yuan et al., 2023; Zhao et al., 2023; Azar et al., 2024; Xu et al., 2024a; Hong et al., 2024; Park et al., 2024; Meng et al., 2024), but comparing their performance in LLM alignment is a complex task that incurs high costs, noise, and the inherent difficulty in judging a response better than another.
 In this work, we provide a comprehensive analysis of PO algorithms, examining their behavior on automatically generated preference datasets.
-We perform this analysis in MuJoCo environments (Todorov et al., 2012), where the underlying reward structure is well defined and offers a clear performance metric to compare agents.
-In particular, we focus on ORPO, as we have found that it largely outperforms DPO in all settings we have considered.
-Our findings indicate that ORPO exhibits distinct failure modes when applied to specific low-quality or noisy datasets.
-These failure modes are present in practical LLM applications, raising concerns about using mixed-quality datasets for PO.
-To find algorithms that are capable of dealing with these more difficult settings, we introduce a novel framework for PO based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO and allows to search for new Preference Optimization (PO) algorithms.
-We perform this search over the space of mirror maps using evolutionary strategies (ES), as they have been shown to be effective for discovering algorithms (Lu et al., 2022; Jackson et al., 2024).
-We outline our contributions in detail below.
+We return to the roots of RLHF by performing this analysis in MuJoCo environments and datasets, where the underlying ground-truth reward structure is well defined and offers a clear performance metric to compare agents.
+In particular, we design a task where a pre-trained agent has to adhere to a new stylistic constraint, emulating the typical conditions of LLM fine-tuning.
+Our findings indicate that many PO algorithms present distinct failure modes when applied to specific mixed-quality or noisy datasets.
+Moreover, we introduce a framework for finding PO algorithms.
+Specifically, we define a class of PO algorithms based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO for particular choices of the mirror map.
+We then show that this class can be easily parametrized and searched using evolutionary strategies (ES), optimizing for the final performance of the trained policy, as measured by the ground truth reward.
+For each setting we consider, we discover an algorithm that significantly outperforms all baselines.
+Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one.
+We use this insight to design a new PO function, Constrained Linear Policy Optimization (CoLPO) that demonstrates promising results in an LLM alingment task.
+We summarize our contributions below.
   1. 1.
-    We perform a systematic analysis of ORPO on preference datasets with varying levels of data quality, noise levels, initial policy, and judge temperature.
+    We perform a systematic evaluation of eight of the most popular PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy.
+We see that most existing algorithms struggle when dealing with noise and mixed-quality date.
   2. 2.
-    We introduce a novel family of offline PO algorithms using mirror descent, which can be easily parameterized and which we explore via ES.
+    We introduce a novel family of offline PO algorithms using mirror descent, named Mirror Preference Optimization (MPO), which can be easily parameterized and explored via ES.
   3. 3.
-    For each failure mode we identify, we find and describe an algorithm within our framework that outperforms ORPO.
-Additionally, we show that allowing losses to vary based on the percentage of training progress drastically boosts performance in some settings.
+    For both noisy and mixed-quality settings, we find and describe a PO algorithm within our framework that largely outperforms all the considered baselines in our MuJoCo benchmark.
   4. 4.
-    We demonstrate that PO algorithms discovered on MuJoCo can generalize to LLM tasks.
-In particular, for a highlighted failure setting, we show that our discovered algorithm improves performance w.r.t.
-the ORPO baseline.
-This finding showcases the efficacy of analyzing PO algorithms in simpler, less computationally expensive environments, as it is possible to obtain insights relevant for LLM applications.
+    We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks.
+In particular, we show that our CoLPO algorithm significantly improves upon the ORPO, CPO, and SLiC-HF baselines.
+Overall, we show that it is possible to gain relevant insights for the design of LLM alignment algorithms through a systematic analysis in simple, computationally inexpensive environments such as MuJoCo.
 2 Preliminaries
-Let $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,T,\mu)$ denote an episodic Markov Decision Process, where $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action spaces, $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ when taking action $a$, $r(s,a)\in[0,1]$ is the reward function, $T$ is the maximum episode length, and $\mu$ is a starting state distribution.
 A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$.
 Let $\tau=\{(s_{t},a_{t})\}_{t=0}^{T-1}$ denote a trajectory of length $T$ and, with a slight overload of notation, let $\pi(\tau)=\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})$ and $r(\tau)=\sum_{t=0}^{T-1}r(s_{t},a_{t})$.
-Lastly, let $\pi(\cdot|\tau)$ be the product distribution of $\pi(\cdot|s_{0}),\dots,\pi(\cdot|s_{N-1})$.
+Lastly, let $\pi(\cdot|\tau)$ be a distribution over $(\Delta(\mathcal{A}))^{T}$ defined as $\pi(\cdot|s_{0})\times\dots\times\pi(\cdot|s_{N-1})$.
+Our objective is to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
+$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)$ (1)
+$:=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
+where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
 Let $\mathcal{D}=\{(s_{0}^{i},\tau_{w}^{i},\tau_{l}^{i})_{i=1}^{N}\}$ be a preference dataset, where each tuple $(s_{0},\tau_{w},\tau_{l})$ consists of a starting state $s_{0}$ and two trajectories with starting state $s_{0}$.
 Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$.
-We assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
-$\mathbb{P}(\tau_{w}\succ\tau_{l})=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$ (1)
+Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
+$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}$ (2)
+$=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$
 where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter.
-Our objective is to exploit the dataset $\mathcal{D}$ to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau):=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})}\sum_{t=0}^{T-1}r(s_{t},a_{t}).$ (2)
-We consider an offline setting, where we do not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
-2.1 Preference optimization
-There are several methods in the literature to optimize the objective in (2) using a preference dataset $\mathcal{D}$.
-We provide here a short review of the main pipelines (Ouyang et al., 2022).
-Two common independent preliminary steps are reward modelling (RM) and supervised fine tuning (SFT).
-RM aims to obtain an estimate of the true reward function, and is usually framed as a maximum likelihood estimation problem for a Bradley-Terry preference model, i.e.
-find
-$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}\log\sigma(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})/\eta),$ (3)
-for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$ and for $\eta\geq 0$.
-SFT is an initial alignment phase, where the starting policy $\pi_{0}$ is trained to imitate high-quality demonstration data.
-In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$, typically doing one epoch over the dataset.
+In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data.
+We also assume the agent does not have access to either the transition probability $P$, the reward function $r$, or the MDP $\mathcal{M}$.
+2.1 Alignment to preference feedback
+There are several algorithms in the literature to optimize the objective in (1) using a preference dataset $\mathcal{D}$.
+We describe supervised fine-tuning (SFT), DPO and ORPO, as they are among the most popular and as many methods can be seen as a variation of one of these algorithms.
+SFT
+SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data.
+In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$.
 We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
-The new objective we want to optimize is
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$ (4)
-where the $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution.
-The expressions in (3) and (4) can be optimized sequentially, first obtaining a reward estimate in (3) and then optimizing (4) with PPO using the reward estimate in place of the true reward.
-Alternatively, the optimization problems in (3) and (4) can be solved implicitly using DPO or ORPO.
-We proceed to discuss each of these methods.
-PPO
-PPO is recognized as one of the preferred methods for optimizing (4) when the necessary computing resources are available, as demonstrated by its success in training state of the art models like GPT-4 (Achiam et al., 2023) and Claude (Antropic, 2023).
-However, it presents a complex pipeline where one needs to effectively train a reward model, perform SFT and then optimize (4), where each phase has a different set of hyper-parameters.
-Additionally, storing the reward model, the reference agent and the current agent in memory is impractical in most setups and often requires sacrificing other aspects, such as batch-size.
-Besides it computational costs, PPO is known to be prone to reward overoptimization (Coste et al., 2024).
 DPO
-Direct Preference Optimization (DPO) bypasses the need for an explicit reward model by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$ (5)
-which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3).
-Refer to Appendix A for details.
-Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs as an alternative to PPO (Yuan et al., 2024; Jiang et al., 2024).
-A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses, which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences.
-In contrast, PPO can leverage the generalization capabilities of the (learned) reward model to generate responses beyond the preference dataset distribution, while the $\mathtt{KL}$-divergence penalty can provide additional regularization.
-To mitigate the risks described above, DPO is usually only applied for a few epochs.
+Direct Preference Optimization (DPO) consists in solving a maximum likelihood estimation problem and a policy optimization problem in a single step.
+The maximum likelihood estimation problem is the one to find an estimate of the true reward function that governs how the preferences are expressed, that is
+$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}g\circ\sigma((r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l}))/\eta),$ (3)
+for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$, for an increasing function $g:\mathbb{R}\leftrightarrow\mathbb{R}$, and for $\eta\geq 0$.
+Here, typically, $g$ is set to be the logarithm function.
+The policy optimization problem is the one to maximize the expected reward, that is
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (4)
+$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$
+where $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution.
+DPO merges these two problems by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
+$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$ (5)
+$\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$
+which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3), with $g=\log$.
+Refer to Appendix C for details.
+Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
+A known issue of DPO is that it pushes probability mass away from the preference dataset and to unseen responses (Xu et al., 2024b), which can cause the final policy to deviate significantly from the reference policy, even when the reference policy aligns well with human preferences.
+To mitigate this risk, DPO is usually applied for a few epochs.
 ORPO
-ORPO is a more recent algorithm that aims to further simplify the training pipeline and, concurrently, to address the distribution shift issue present in DPO.
-ORPO merges the SFT and DPO steps into one, which optimizes the unified objective
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\right]$ (6)
+ORPO further simplifies the training pipeline and addresses the distribution shift issue present in DPO. It merges the SFT and DPO steps into one, optimizing the unified objective
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg{[}\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}$ (6)
+$+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg{]}$
 where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$.
-ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, which substitutes the role of the SFT step and of the reference model in preventing the optimized policy from moving too far away from the dataset distribution.
+ORPO gets rid of the need for a reference model by adding an SFT term to the preference optimization objective function, and uses this term to prevent the optimized policy from moving too far away from the dataset distribution.
 Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.
+Research on preference optimization has been very active and many methods have been proposed. We present a summary of the most popular algorithms in Table 1 and a brief discussion in Appendix A.
+Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm.
+Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement.
+For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix F.
+Method Objective
+RRHF (Yuan et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max\left(0,-\frac{1}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)+\frac{1}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)\right)$
+SLiC-HF (Zhao et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max(0,\delta-\log\pi_{\theta}(\tau_{w}|x)+\log\pi_{\theta}(\tau_{l}|x))$
+DPO (Rafailov et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}\right)$
+IPO (Azar et al., 2024) $-\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
+CPO (Xu et al., 2024a) $\log\pi_{\theta}(\tau_{w}|x)+\log\sigma(\beta\log\pi_{\theta}(\tau_{w}|x)-\beta\log\pi_{\theta}(\tau_{l}|x))$
+ORPO (Hong et al., 2024) $\log\pi(\tau_{w})+\lambda\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)$
+R-DPO (Park et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}+(\alpha|\tau_{w}|-\alpha|\tau_{l}|)\right)$
+SimPO (Meng et al., 2024) $\log\sigma\left(\frac{\beta}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)-\frac{\beta}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)-\gamma\right)$
+Table 1: A summary of popular PO algorithms. The objective in each row is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
 2.2 Mirror Maps
@@
 The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
-Definition 2.1 ($\omega$-potential mirror map Krichene et al. (2015)).
+Definition 2.1 ($\omega$-potential mirror map (Krichene et al., 2015)).
 For $u\in(-\infty,+\infty]$, $\omega\leq 0$, an $\omega$-potential is defined as an increasing $C^{1}$-diffeomorphism $\phi:(-\infty,u)\rightarrow(\omega,+\infty)$ such that
@@
 $h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
-When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix B).
+When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix D).
 Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one.
@@
 An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
-$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s)):=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
+$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))$
+$-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
 where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$.
-When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix B).
+When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix D).
 When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
-3 Methodology
-We develop a new framework for preference optimization based on mirror maps, which generalizes DPO and ORPO.
-We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a more general Bregman divergence, that is, we aim to solve the problem
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$ (7)
+2.3 Evolution Strategies
+OpenAI-ES  (Salimans et al., 2017) is a popular method to be able to optimize non-differentiable functions and it has been widely used to meta-learn objectives  (Lu et al., 2022; Jackson et al., 2024), as it obtains an unbiased estimate of the gradient (unlike second order gradient methods).
+The gradient $\nabla_{\zeta}F(\zeta)$ is estimated using:
+$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$
+where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
+3 Mirror Preference Optimization
+We introduce Mirror Preference Optimization (MPO), a new framework for preference optimization that generalizes DPO and ORPO.
+We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a Bregman divergence and aim to solve the problem
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (7)
+$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$
 where $D_{h}$ is the Bregman divergence induced by a mirror map $h$.
@@
 where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
-We provide a proof for Theorem 3.1 in Appendix A. By plugging (8) in the maximum likelihood problem in (3), we obtain the objective:
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}\log\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$ (9)
+We provide a proof for Theorem 3.1 in Appendix C. By plugging (11) in the maximum likelihood problem in (3), we obtain the objective:
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}g\circ\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))$ (9)
+$-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$
 where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$.
-When $\phi=e^{x}$, the expression in (7) recovers the DPO objective in (5).
-Additionally, we can modify the expression in (7) to include an SFT term in order to avoid the SFT step, in a manner similar to ORPO.
-Let $\psi$ be an $\omega$-potential and let $\pi_{\mathrm{ref}}$ be the uniform distribution, then the new objective we consider is
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\psi(\pi(\tau_{w}))+\lambda\log\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\right],$ (10)
-where the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ do not appear as they have cancelled out.
+Two-step MPO
+We can use the objective in (9) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (9).
+When $\phi=e^{x}$, the (9) is equivalent to (5) and we recover the DPO algorithm.
+One-step MPO
+By adding an SFT term to (9) and by setting the $\pi_{\mathrm{ref}}$ be the uniform distribution, we obtain a class of one-step PO algorithms.
+These algorithms consists in a single phase, where we optimize the objective
+$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\psi(\pi(\tau_{w}))$ (10)
+$+\lambda g\circ\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big{]},$
+where $\psi$ is an $\omega$-potential.
+Here, the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ have canceled out due to $\pi_{\mathrm{ref}}$ being uniform.
 We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization.
-When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x)-\log(1-x)$, (10) recovers the ORPO objective in (6).
+When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (10) recovers the ORPO objective in (6).
 The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification.
-In the following, we will show that taking into account the properties of the preference dataset when choosing $\psi$ and $\phi^{-1}$ can lead to a better performance of the trained policy.
-Following the intuition developed by Jackson et al.
-(2024), we also allow our objective function, specifically $\psi$ and $\phi^{-1}$, to account for training progress.
-This is equivalent to changing the Bregman divergence penalty in (7) during training and, as we shall discuss in the following section, helps in dealing with datasets with mixed-quality data.
-3.1 Learning mirror maps
-To search the space of PO algorithms we have defined, we employ a neural network parametrization for both $\psi$ and $\phi^{-1}$, which we optimize using evolutionary strategies.
-Similarly to Alfano et al.
-(2024), we parameterize both $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
-$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),\;e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
+In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform current baselines.
+3.1 Meta Learning PO objectives
+To search the space of PO algorithms we have defined, we employ a neural network parametrization for $g$, $\psi$, and $\phi^{-1}$, which we optimize using evolutionary strategies (Salimans et al., 2017).
+Similarly to (Alfano et al., 2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
+$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),$
+$e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
 where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$.
-The non-negative kernels and the increasing activation functions guarantee the monotonicity of $\psi$ and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions.
-To ensure that we are able to recover the ORPO objective, we add $a\log(x)$ and $b\log(x)-b\log(1-x)$ to the final outputs of $\psi$ and $\phi^{-1}$, respectively, where $a,b\geq 0$.
-In case we want to take into account training progress, we give a second input to the neural network, i.e.
-$x\cdot n/N$, where $n$ is the current epoch and $N$ is the total number of epochs through the dataset.
-To ensure that monotonicity is preserved, we lower bound the weights associated to the second input with the negative of the respective weights associated to the first input.
-To search for the best $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy (Salimans et al., 2017).
-Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$.
+The non-negative kernels and the increasing activation functions guarantee the monotonicity of $g$, $\psi$, and $\phi^{-1}$, while the several different activation functions facilitate expressing complex functions.
+To ensure that we are able to recover the DPO and ORPO objectives, we add $a\log(x)$, $b\log(x)$ and $c\log(x/(1-x))$ to the final outputs of $g$, $\psi$ and $\phi^{-1}$, respectively, where $a,b,c\geq 0$.
+To search for the best $g$, $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy.
+Denote by $\zeta$ the parameters of $g$, $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$.
 Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e.
 $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$.
-We estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
-$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$
-where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
 We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient.
-4 Experiments
-We carry out our experiments on continuous reinforcement learning tasks in MuJoCo and on LLM fine-tuning.
-In particular, due to computational efficiency constraints, we discover mirror maps in MuJoCo and report their performance when evaluated for an LLM finetuning task.
-4.1 Environments
-We consider three different preference optimization settings which we describe below.
-Each setting includes a starting policy, a set of preference datasets with different properties and an evaluation method for the trained policy.
-We apply ES to the first two settings and test the discovered objectives on the last one.
-Hopper
-Our objective in Hopper is to learn a policy from scratch, i.e.
-with a randomly initialised policy, using a preference dataset.
-On Hopper, we train four reference agents of different skill levels, specifically, of respective expected cumulative reward (or value) of 900, 1200, 1800, and 2100 (the expert agent).
-Each dataset consists of 5120 rows, each with two trajectories of length 1000 starting from the same state.
-We generate a dataset by comparing trajectories by the expert agent (average return 2100), with other trajectories by one of our reference agents.
-A Bradley-Terry judge ranks each pair of trajectories, based on their true reward.
-The trained policy is then evaluated on the standard Hopper environment.
-To test the efficacy and robustness of our loss functions, we analyse a few variations of these standard datasets that represent common issues of real world data.
-  • Base dataset: we compare a trajectory from the expert agent with one of a reference agent.
-  • Noisy dataset: we flip a varying portion of the rankings in the base dataset.
-  • Shuffled dataset: the comparisons are not always between the expert agent and a second agent.
-Namely, $25\%$ of the comparisons are between two trajectories from the expert agent, $50\%$ of the comparisons are between a trajectory of the expert agent and one of a reference agent, and $25\%$ of the comparisons are between two trajectories of a reference agent.
-  • Poor judge dataset: the judge is more likely to flip labels when the two trajectories have closer reward values. This is implemented as an increase in the temperature of the Bradley-Terry judge.
-Ant
-To further investigate a scenario similar to fine-tuning a large language model (LLM), we examine a setting in which a pre-trained agent is required to modify its behavior to achieve its goal with a stylistic constraint.
-Specifically, in the Ant environment, we take an agent that has been pre-trained on the standard Ant goal of moving forward, and enforce the objective of avoiding the use of one of its legs.
+In practice, to compute (12), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions.
+We then train 256 agents with the perturbed objective functions on an a preference dataset.
+To measure the value of each agent, i.e.
+$\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent.
+Refer to Appendix E for further discussion and details on the ES methodology.
+In our experiments, we consider both the case where we fix $g=\log$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions.
+We also perform the evolution on both the two-step and the one-step MPO classes.
+Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
+RRHF 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
+SLiC-HF 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
+DPO 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
+IPO 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
+CPO 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
+ORPO 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
+R-DPO 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
+SimPO 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
+SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
+LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
+With $g=\log$
+1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
+1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
+2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
+2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
+With parametrized $g$
+1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
+2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
+1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
+2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
+1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
+Table 2: Three Legged Ant (TLA) Adaptation.
+Performance of popular and discovered MPO algorithms on TLA, for various dataset settings.
+For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents.
+For discovered MPO algorithms, we report performance of one-step MPO (1S-MPO) and two-step MPO (2S-MPO).
+For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters).
+We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
+4 MuJoCo Experiments
+Our first set of experiments is carried out on continuous RL tasks in MuJoCo.
+In particular, we show the performance of all the algorithms presented in Table 1 across several settings and we compare it with the performance of our discovered objectives.
+To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries.
+We report our hyper-parameters in Appendix I.
+4.1 Tasks
+To reproduce the typical conditions of LLM fine-tuning, which involve a pre-trained model, we consider a setting where the task is to adapt a pre-trained agent to meet the original objective while adhering to an additional stylistic constraint.
+Specifically, in the Ant environment, we start from an agent that has been pre-trained on the standard Ant goal of moving forward and enforce the objective of avoiding the use of one of its legs.
 This is accomplished by introducing the Three-legged-ant (TLA) environment, a modified version of Ant where utilizing the fourth leg results in significant penalties.
-We train one agent in the original Ant environment, achieving a reward of 6000, and another in the TLA environment, which achieves a reward of 3900.
-For comparison, the agent trained in the original Ant environment achieves a reward of 1700 when evaluated in the TLA environment.
-The dataset generation follows the same protocol as described for the Hopper environment, where trajectories are collected from both the Ant and TLA agents.
-However, in this setting the number of rows for each dataset is 1280, to account for the fact that we do not want the agent to learn from scratch, but to adjust its policy to a new instruction.
-The trained policy is then evaluated on the TLA environment.
-LLM tuning
-Finally, we evaluate one of our discovered objective functions on a real-world LLM fine-tuning task.
-To simulate a scenario involving mixed data quality, we fine-tune the gemma-7b [^2] https://huggingface.co/google/gemma-7b model on a modified version of the dpo-mix-7k [^3] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset, where half of the responses, selected at random, are replaced with responses generated by gemma-2b [^4] https://huggingface.co/google/gemma-2b, a model that typically produces lower-quality responses compared to those originally present in DPO-mix.
-This approach replicates the shuffled dataset described for the MuJoCo experiments, aiming to simulate the challenges of training with datasets of varying quality, a common issue in fine-tuning tasks where it is difficult and costly to ensure uniformly high-quality data.
-We implement our experiments in MuJoCo using the brax library (Freeman et al., 2021) for the environments and the evosax library (Lange, 2022) for the evolution.
-We use the same hyper-parameters for all MuJoCo datasets, which have been tuned for ORPO using Weights and Biases [^5] https://wandb.ai/site.
-We report them in Appendix C, along the hyper-parameters for OpenES.
-For the LLM tuning, we modify the Alignment Handbook library (Tunstall et al., ), and use the default hyper-parameters.
-4.2 Results: Mujoco
-Table 1: Results for Hopper.
-For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset.
-When ranking two trajectories of value 2100 and 1800, a judge with low, medium, and high temperature provides the correct ranking 95%, 85% and 75% of the time, respectively.
-Value agent 1 Value agent 2 Noise Shuffled Judge temp. ORPO Discovered
-2100 900 0 No Low 2003$\pm$20 2012$\pm$11
-2100 1200 0 No Low 2055$\pm$12 2055$\pm$10
-2100 1800 0 No Low 2043$\pm$ 15 2073$\pm$12
-2100 2100 0 No Low 2070$\pm$12 2098$\pm$11
-2100 900 0.1 No Low 1519$\pm$45 1724$\pm$43
-2100 1800 0.1 No Low 1936$\pm$25 1945$\pm$25
-2100 2100 0.1 No Low 2112$\pm$9 2094$\pm$13
-2100 900 0.2 No Low 662$\pm$49 1127$\pm$45
-2100 900 0.3 No Low 623$\pm$29 698$\pm$26
-2100 1800 0 No Medium 1884$\pm$22 1928$\pm$23
-2100 1800 0 No High 1858$\pm$32 1895$\pm$14
-2100 900 0 Yes Low 975$\pm$51 1099$\pm$35
-Table 2: Results for TLA. For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the objective discovered for that dataset.
-Value agent 1 Value agent 2 Noise Shuffled ORPO Discovered
-3900 1700 0 No 3277$\pm$49 3473$\pm$63
-3900 1700 0 Yes 2425$\pm$58 3485$\pm$163
-3900 1700 0.1 No 2675$\pm$62 3837$\pm$52
-Table 3: Results for temporally-aware objectives.
-For each dataset, we report the average value and standard error of 25 agents trained using ORPO and the temporally-aware objective discovered for that dataset.
-Value agent 1 Value agent 2 Env Noise Shuffled ORPO Discovered
-2100 900 Hopper 0 Yes 975$\pm$51 2032$\pm$13
-3900 1700 TLA 0 No 3277$\pm$49 3903$\pm$42
-3900 1700 TLA 0 Yes 2425$\pm$58 3251$\pm$83
-We provide the results of our experiments for MuJoCo in Tables 12, and 3, which report the performance of ORPO and our discovered objectives for Hopper, TLA, and the temporally aware case, respectively.
-To provide information on the behavior of these objectives as well as guidelines on how to design loss landscapes for the settings we have considered, we compare the landscapes of ORPO and of the discovered objectives in Figures 123, and 4.
-In these figures, we report, in absolute value, the gradients of the objective with respect to $\log(\pi(\tau_{w}))$ and $\log(\pi(\tau_{l}))$, as well as a sampled training trajectory.
-We discuss the results below, by dataset type.
-Base Dataset
-For the base dataset, we observe different behaviors for the Hopper and TLA environments.
-For the Hopper dataset, we see that ORPO achieves the performance of the best reference agent in all cases and that it is not significantly outperformed by the learned objective.
-On the other hand, ORPO does not reach the performance of the best reference agent in the TLA dataset and is surpassed by the discovered objectives.
-Remarkably, the temporally aware objective achieves the performance of the best reference agent, presenting a $19\%$ improvement over ORPO.
-As shown by Figures 1 and  4, both the static and the temporally aware discovered objectives present a smaller gradient than ORPO, which means that they cause smaller updates.
-Additionally, for ORPO we have that if $\log(\pi(\tau_{w}))<\log(\pi(\tau_{l}))$ then the gradients increases suddenly by half, while for the temporally aware discovered objective we can observe that the gradient increases more slowly.
-Essentially, this means that ORPO causes a large update when the agent mistakenly assigns more probability to the rejected trajectory, while the temporally aware discovered objective causes a large update only if the agent assigns significantly more probability to the rejected trajectory.
-Figure 1: The plots represent the gradients, in absolute value, of ORPO (left) and the objective discovered on the base TLA dataset (right). The dots represent a sampled training trajectory.
-Noisy Dataset
-As expected, the performance of ORPO declines when the dataset contains randomly flipped labels, with the degradation becoming more pronounced as the percentage of flipped labels increases.
-While the performance degrades for our discovered objectives in Hopper as well, it happens more slowly and we show significant improvements over ORPO for all noise levels.
-In TLA, the discovered objective is capable of almost achieving the best reference agent value despite the noise.
-Figure 2 shows that the objective discovered in Hopper assigns a roughly constant value to the gradient w.r.t.
-the chosen logits, while the gradient w.r.t.
-the rejected logits increases as $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ decreases.
-Consequently, when the agent encounters a flipped label where $\log(\pi(\tau_{w}))\ll\log(\pi(\tau_{l}))$, the agent does not take larger updates to increase the chosen logits, but only larger updates to decrease the rejected logit, thus mitigating the noise effects.
-On the other hand, the objective discovered in TLA manages to mitigate the influence of noise by reducing the gradient value and by taking larger updates only when the chosen logits are much larger than the rejected logits.
-Figure 2: The plots represent the gradients, in absolute value, of the objectives discovered on the noisy Hopper dataset (left) and on the noisy TLA dataset (right).
-The level of noise is 0.1 for both plots.
-The dots represent a sampled training trajectory.
-Shuffled Dataset
-While effectively using the same data as the base dataset, the performance of ORPO significantly decreases in both the Hopper and TLA shuffled dataset, where low quality trajectories are selected as the preferred choice when compared to other low quality trajectories.
-In contrast, the discovered objectives achieve a value of the final policy close to the best reference agent.
-In particular, the value reached by the temporally aware objective for Hopper is more than double the value reached by ORPO.
-All discovered objectives present smaller gradients than ORPO, and therefore induce smaller updates, which prevents taking large updates when the chosen logits belong to a low-quality trajectory.
-As to the temporally aware objective discovered on HopperFigure 4 shows that it induces very small updates for the chosen logits when they are low, and therefore probably belong to a low-quality trajectory.
-Lastly, we note a decrease in the absolute value of the gradient at the top and right borders as training progresses, which induce smaller updates.
-Figure 3: The plots represent the gradients, in absolute value, of the objectives discovered on the shuffled Hopper dataset (left) and on the shuffled TLA dataset (right).
-The dots represent a sampled training trajectory.
-Poor judge
-As expected, we observe a decrease in performance for ORPO when we increase the temperature of the judge.
-However, since we compared trajectories from two high-quality reference agents, the drop in performance is small.
-This is also reflected in the discovered objectives, which do not manage to significantly outperform ORPO.
-Figure 4: The plots represent the gradients, in absolute value, of the temporally aware objectives discovered on the shuffled Hopper dataset (left) and on the base TLA dataset (right), for three different values of the training progress.
-For each row, the dots represent the respective third of a sampled training trajectory.
-4.3 Results: Transfer
-We show that PO objectives learnt with our methodology can transfer to other domains.
-In particular, the static objective discovered on the shuffled TLA dataset tranfers to both the shuffled Hopper dataset, where it obtains a final policy value of 1735$\pm$39, and to the LLM-tuning task with shuffled data defined above.
-We train the base LLM on the modified DPO-mix dataset with both ORPO and the discovered objective.
-We observe that the first methodology achieves 57% accuracy on the test set while the latter achieves 62%.
-Additionally, we compare the two trained models with AlpacaEval (Li et al., 2023), and obtain a 53% winrate for the model trained with the discovered objective.
-5 Related Work
+The offline preference optimization task is defined as follows.
+We train one agent (the original agent) in the original Ant environment, achieving a reward of 6000, and another (the target agent) in the TLA environment, which achieves a reward of 3900.
+For comparison, the original agent achieves a reward of 1700 in the TLA environment.
+We then generate a preference dataset of 1280 rows, each with two trajectories of length 1000 starting from the same state.
+Each trajectory is generated by either the original or the target agent, depending on the current setting.
+A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward.
+We also consider training a randomly initialized agent on the Hopper environment, using a preference dataset with 5120 pairs of trajectories.
+This setting addresses the case where a new behavior has to be learned completely from the preference dataset and no prior knowledge of the task is available.
+Due to space constraints, we report the results of the experiments for this task in Appendix H.
+We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
+  • Base dataset: for each pair of trajectories, one is generated by the original agent and one by the target agent.
+  • Noisy dataset: same as the base dataset but each label given by the judge is flipped with probability $\varepsilon$.
+  • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one.
+The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between a trajectory of the target agent and one of the original agent, and $25\%$ comparisons between two trajectories of the original agent.
+Figure 1: SimPO
+Figure 2: Discovered 1S-MPO (shuffled)
+Figure 3: ORPO
+Figure 4: Discovered 1S-MPO (noisy)
+Figure 5: Plots of the absolute value of the gradient of SimPO, ORPO, and of two discovered objectives, i.e.
+1S-MPO (shuffled) and 1S-MPO (noisy, $\epsilon=0.3$).
+The dots are sampled datapoints from the training distribution, with lighter points representing more recent datapoints during the course of training.
+In the bottom right and corner of each plot, SimPO’s and ORPO’s gradients go to zero, while the gradient of our discovered loss functions keep increasing.
+The effect of this on training is evident observing the datapoints: in ORPO and SimPO, the likelihood of the rejected points is not minimized, leading to decreased performance.
+4.2 Results
+We provide the results of our experiments for TLA in  Table 2, which reports the performance of several PO algorithm and of our discovered objectives.
+We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters.
+All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
+First of all, we notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a large drop in performance in the mixed-quality and noisy settings.
+Importance of SFT
+The first group within Table 2 shows that SFT plays a key role in the performance across different settings.
+SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings.
+IPO and DPO, which do not contain an SFT term but have an SFT step, are among the top performers on the base, mixed-quality and low noise settings.
+Their performance finally drops when the noise level reaches 0.3.
+Lastly, the algorithms that present an SFT term in their objectives, e.g.
+CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
+Keep optimizing
+Furthermore, while RRHF and SliC-HF are very similar, SliC-HF allows the PO part of the objective to be clipped when $\pi(\tau_{w})>\pi(\tau_{l})+\delta$, rather than when $\pi(\tau_{w})>\pi(\tau_{l})$.
+This modification leads to a higher performance on all tasks, demonstrating that it is beneficial to keep optimizing the policy even if $\pi(\tau_{w})>\pi(\tau_{l})$.
+To further stress this point, we consider the objective
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big{]},$
+which we call Linear Preference Optimization (LPO).
+LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table 2, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
+Discovered objectives
+Table 2 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it.
+We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings.
+Differently from the human-designed algorithms, the discovered objectives recover the performance of the target agent in multiple instances and are more robust to the mixed-quality and noisy settings.
+We note that allowing the evolution procedure to learn $g$ leads to a better performance in all dataset settings.
+Additionally, we have that the objectives discovered within the two-step MPO class always have a lower performance than those within the one-step MPO class.
+This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
+When exploring the one-step MPO class with $g=\log$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
+$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\alpha\log(\pi(\tau_{w}))$
+$+\lambda\log\sigma\left(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\right)\big{]},$
+where the coefficients $\alpha$ and $\beta$ depend on the setting.
+In particular, we have a low value for $\alpha$ and a high value for $\beta$ in the noisy settings, while we observe the opposite in the mixed-quality setting.
+These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories.
+When we search the two-step MPO class with $g=\log$, we recover DPO.
+Figure 5 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$.
+In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings.
+For comparison, we provide the same plots for the ORPO and SimPO objectives.
+The hand-crafted algorithms present a larger gradient when $\pi(\tau_{w})<\pi(\tau_{l})$ and a smaller one when $\pi(\tau_{w})>\pi(\tau_{l})$, that is, they induce large updates when the data-point contradicts the current behaviour of the agent, and small otherwise.
+The objective discovered on the noisy dataset has the opposite behavior, meaning that it only optimizes the more robust-to-noise SFT term when $\pi(\tau_{w})<\pi(\tau_{l})$.
+As $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ becomes larger, it shifts toward increasingly large updates thanks to the PO term.
+A similar pattern appears in the mixed-quality dataset, where the objective also increases its updates as the difference $\log(\pi(\tau_{w}))-\log(\pi(\tau_{l}))$ grows.
+The key distinction between these two losses is that the shuffled loss triggers high-gradient updates even when $\pi(\tau_{w})<\pi(\tau_{l})$.
+We highlight once more that both discovered objectives advocate to keep optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
+For additional results and analysis for the Hopper task, we refer the reader to Appendix H.
+WR $s_{\text{WR}}$ LCWR $s_{\text{LCWR}}$ avg length
+SLiC-HF 16.77 1.32 13.18 0.28 3016
+ORPO 24.47 1.52 19.25 0.29 1679
+CPO 25.96 1.54 21.07 0.25 1623
+CoLPO 31.94 1.65 24.01 0.21 2391
+Table 3: Alpaca Eval LLM results.
+We report Win Rate (WR), standard error for Win Rate ($s_{\text{WR}}$), Length Controlled Win Rate (LCWR), standard error for LCWR ($s_{\text{LCWR}}$) and average length of responses.
+We trained gemma-7b on dpo-mix-7k for 3 epochs.
+We then compared our model to GPT-4 using Llama-3.3-70B-Instruct as judge.
+All algorithms use the same $\beta=0.5$ and the same learning rate of $55e-$ 5.
+In the case of CoLPO and SLiC-HF, we used the same cutoff rate at 5.
+5 Experiments: LLM transfer
+We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting.
+In particular, we found that all our discovered loss functions kept optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, while all algorithms in Table 1 have a PO term whose gradient quickly reaches 0 when $\pi(\tau_{w})>\pi(\tau_{l})$.
+Using this insight, we design a new PO objective, which we call Constrained Linear Preference Optimization (CoLPO) and define as
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(\tau_{w}|x)$
+$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)<\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),0)$
+$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)>\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),\gamma)\big{]},$
+where $\beta\geq 0$, $\gamma\geq 0$ and $\alpha\leq 0$ are hyperparameters.
+The CoLPO objective keeps optimizing the policy when $\pi(\tau_{w})>\pi(\tau_{l})$, but only if $\log\pi_{\theta}(\tau_{w}|x)$ is over a certain threshold.
+This constraint is placed to stop the optimizer from increasing the difference $\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))$ by descreasing both probabilities.
+We evaluate CoLPO on an LLM fine-tuning task, where we fine-tune the gemma-7b (Team et al., 2024) model on the dpo-mix-7k [^2] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset for 3 epochs.
+Being a real-world preference dataset, dpo-mix-7k naturally presents characteristics of the mixed-quality and the noisy settings.
+To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the CoLPO objective.
+We evaluate the tuned LLM against GPT-4, using the AlpacaEval library  (Li et al., 2023).
+All models were trained using the default hyper-parameters, with $\beta=0.5$, $\gamma=5$ (for CoLPO and SLiC-HF), and $\alpha=-1$ (for CoLPO).
+For details on hyperparameters and training settings consult Appendix J.
+Table 3 reports the results of CoLPO, as well as the results for other one-step algorithms, and shows that CoLPO largely outperforms existing algorithms, confirming the validity of the insights gained in the MuJoCo simulations.
+Lastly, we report that the the log-probabilities for the final training batch of chosen and rejected responses are -1.17 and -11.38 respectively for CoLPO, compared to -0.61 and -2.98 for ORPO, respectively.
+This results further confirms that CoLPO is effective in decreasing the probability of rejected responses and mantaining high the probability of chosen responses.
+6 Conclusion
+We have introduced a novel framework for Preference Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolutionary strategies.
+Through a systematic evaluation across diverse settings in MuJoCo environments, we have demonstrated that the performance of our discovered objectives consistently exceeds the performance of existing methods, particularly in noisy and mixed-quality datasets where many baselines underperform.
+Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$.
+To address this, we proposed a simple objective, CoLPO, that avoids such loss truncation while preventing the optimizer from decreasing the probability of chosen datapoints.
+We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
+Appendix A Popular PO Algorithms
+These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT.
+The comparison also includes IPO, a theoretically-based approach that handles pairwise preferences differently than DPO.
+Another method, CPO, combines sequence likelihood as a reward with an SFT objective.
+Finally, R-DPO modifies the original DPO by adding regularization to prevent length exploitation.
+Appendix B Related Work
 Automatic Discovery of Preference Optimization Loss Functions
 Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024).
-An approach particularly relevant to our method is DiscoPOP by Lu et al.
-(2024), which leverages an LLM to discover objective functions for LLM tuning.
-They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by Tang et al.
-(2024).
+An approach particularly relevant to our method is DiscoPOP by (Lu et al., 2024), which leverages an LLM to discover objective functions for LLM tuning.
+They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by (Tang et al., 2024).
 Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space.
@@
 Generalisations of DPO
-A generalization of DPO alternative to ours is $f$-DPO, developed by Wang et al.
-(2023), which consists in replacing the $\mathtt{KL}$-divergence in (2) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function.
+A generalization of DPO alternative to ours is $f$-DPO, developed by (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function.
 We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa.
 They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms.
-Huang et al.
-(2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
-6 Conclusion
-We have introduced a novel framework for Policy Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolution strategies.
-Through a systematic evaluation across diverse settings in MuJoCo environments, we demonstrated that our discovered objective functions consistently match or exceed the performance of existing methods, particularly in noisy and mixed-quality datasets where the ORPO baseline struggles.
-The introduction of temporally aware objective functions further improved performance, allowing the optimization process to vary and adapt during training.
-Analysing the landscape of the discovered objectives, we give an intuition that justifies improved performance as well as guidance in the design of new PO algorithms.
-Our results also indicate that the discovered objectives generalize beyond simple reinforcement learning environments, showing promising performance when transferred to LLMs, thereby confirming the broader applicability of our approach.
-Appendix A Proof of Theorem 3.1
-We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follow by taking $\phi=e^{x}$
-Theorem A.1 (Theorem 3.1).
+(Huang et al., 2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
+Appendix C Proof of Theorem 3.1
+We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
+Theorem C.1 (Theorem 3.1).
 Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
@@
 The theorem statement is obtained by rearranging the last equation and denoting $c(s_{0})=\lambda$
-Appendix B Further discussion of $\omega$-potentials
+Appendix D Further discussion of $\omega$-potentials
 We show here two examples of Bregman divergence induced by an $\omega$-potential mirror map, that is when $\phi(x)=e^{x-1}$ and when $\phi(x)=x$.
@@
 which is the definition of the Euclidean distance.
-Appendix C Hyper-parameters
-We give the hyper-parameters we use for training in Tables 4 and 5.
+Appendix E Further discussion on Evolution Strategies
+Evolution Strategies (ES) represent a powerful, backpropagation-free method for optimizing complex functions, that has been particularly successful in the context of long-horizon, noisy, and bi-level optimization tasks such as RL and meta-RL.
+ES, and in particular the OpenAI-ES algorithm (Salimans et al., 2017), rely on perturbation-based sampling to estimate gradients without requiring backpropagation through the entire computational graph.
+This feature makes ES well-suited for tasks with long computational graphs, for instance algorithms with many updates, where, due to memory constraints, traditional gradient-based methods have to resort to gradient truncation, introducing bias (Werbos, 1990; Metz et al., 2022; Liu et al., 2022).
+In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section 3.1, so that an agent trained using the objective in (10) achieves the highest value.
+Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$.
+Lastly, let $F(\zeta)$ be the expected cumulative reward (or value) of $\pi^{\zeta}$, i.e.
+$F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$.
+At each iteration, we estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
+$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$ (12)
+where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
+Appendix F Further discussion on Online vs Offline Methods
+In the domain of RL and preference optimization, the choice between online and offline algorithms presents a critical trade-off, influencing computational efficiency, data requirements, and generalization capabilities.
+Online methods, such as PPO, iteratively collect and incorporate new data during training.
+These inherently support exploration of the environment, enabling the discovery of novel strategies or behaviors that are not captured in pre-existing datasets.
+However, they need feedback for each generated “trajectory” (or response, in the LLM case), which might be expensive to obtain.
+Online methods are also more complex and particularly sensitive to hyperparameters, often requiring meticulous tuning for stability and efficiency.
+Offline algorithms, such as DPO and its variants, rely entirely on pre-collected datasets.
+These methods are designed for efficiency and simplicity: they don’t require any additional feedback from users and are therefore particularly effective in scenarios where feedback is delayed or unavailable.
+However, the reliance on static datasets means offline methods may struggle to generalize beyond the training data, particularly if the distribution shift between the training dataset and test time distribution is significant.
+Additionally, the performance of the algorithm is closely tied to the quality of the training dataset: noisy, biased, or corrupt datasets can severely degrade performance, as these methods cannot mitigate such issues through exploration or resampling.
+In summary, RLHF (i.e., online) is considered the superior approach, particularly when substantial amounts of online labels are accessible.
+This makes it the industry standard (Xu et al., 2024b).
+While DPO has been theoretically equated to optimizing using PPO and a reward model trained on an offline dataset, recent empirical research (Tang et al., 2024) has challenged this notion.
+These studies have demonstrated that online methods, such as PPO, consistently outperform offline methods like DPO.
+This superiority is attributed to the benefits of on-policy sampling.
+While DPO has occasionally outperformed PPO, it’s important to note that several studies (Xu et al., 2024b; Song et al., 2024) have consistently shown PPO’s overall superiority.
+DPO’s relative strength lies in its simpler training regime, which avoids the complexities associated with reward model inaccuracies.
+However, DPO’s performance is significantly limited by its sensitivity to distribution shift, especially when the offline preference data lacks diversity (Song et al., 2024).
+This limitation becomes particularly evident when querying the model with out-of-distribution data, a common challenge for methods relying solely on offline data.
+To mitigate this issue, DPO-iter (Xu et al., 2024b), which incorporates online data, has been proposed as a potential solution.
+Appendix G Further discussion on Meta-Learning Algorithms
+Meta-learning, or “learning to learn”, has been extensively employed to automate the design of algorithms that can either adapt rapidly with minimal data samples or generalize effectively to unseen data, tasks, or environments.
+The development of broadly applicable algorithms is particularly critical in the context of preference optimization for LLMs.
+Here, LLMs are fine-tuned on relatively small datasets of offline data but must generalize to a virtually infinite range of potential user queries.
+Prior work in meta-learning has demonstrated success in developing generalizable optimization algorithms and loss functions (Lu et al., 2022; Jackson et al., 2024; Lu et al., 2024; Goldie et al., 2024; Kirsch et al., 2020).
+At its core, meta-learning is defined as a bilevel optimization problem with an inner and an outer loop.
+The inner loop consists in an iterative optimization algorithm that trains agents to solve a predetermined task given a set of meta-parameters.
+The outer loop consists in evaluating the agents trained in the inner loop and update the meta-parameters accordingly, following some optimization method like second order gradient descent (Finn et al., 2017).
+The evaluation of the agents is typically done on a held-out dataset in supervised learning or by sampling trajectories on the environment simulator in RL (Lu et al., 2022; Jackson et al., 2024).
+In our setting, the inner loop is the offline preference optimization algorithm, while the outer loop is the agent evaluation on the environment (online) and the update of the meta-parameters $\zeta$.
+Base Mixed Quality Noisy $\epsilon=0.1$
+DPO 1796 $\pm$ 78 458 $\pm$ 82 693 $\pm$ 131
+IPO 2049 $\pm$ 13 1606 $\pm$ 91 739 $\pm$ 118
+CPO 2078 $\pm$ 12 1078 $\pm$ 35 1813 $\pm$ 33
+ORPO 2022 $\pm$ 15 1039 $\pm$ 20 1710 $\pm$ 33
+SimPO 2027 $\pm$ 15 1460 $\pm$ 94 1794 $\pm$ 65
+Table 4: Hopper. Perfomance of popular MPO algorithms on the Hopper setting. The agent is randomly initialised.
+Appendix H MuJoCo Additional Results
+Below, we report a simplified version of the discovered loss on the shuffled and noisy $\epsilon=0.3$ datasets:
+$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (13)
+$+1.4\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
+$+0.22\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2}$
+$+0.2\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{3}$
+$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (14)
+$+\max((1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2},$
+$0.1\cdot(1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))))$
+We consider a second set of simulations on MuJoCo, based on the Hopper environment.
+As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries.
+We report our hyper-parameters in Appendix I.
+H.1 Hopper Tasks
+Differently from the TLA task, we consider a setting where the agent is randomly initialized and needs to learn a policy from scratch.
+On Hopper, we train an agent with an expected cumulative reward of 2100 (the expert agent) and an agent with an expected cumulative reward of 900 (the bad agent).
+We generate the preference datasets in the same way we do for TLA, with the exceptions that the number of rows is 5120 and the trajectories are generated by either the expert or the bad agent.
+We consider the same three variations of the preference datasets used in TLA, where the expert agent corresponds to the target agent, and the bad agent corresponds to the original agent.
+We include more data compared to the TLA setting as it takes more datapoints for the agent to learn from scratch rather than to adapt to a slightly different objective (like in the TLA case).
+H.2 Further MuJoCo Analsys
+In addition to the noisy dataset, we also considered a bad judge setting, where the judge would be more likely to swap the label of a pair of trajectories if their ground truth rewards were closer to each other.
+This is practically implemented as an increase in the temperature of the Bradley-Terry judge.
+However, we did not notice significantly different results compared to the simple noisy setting, therefore detailed results are not reported.
+H.3 Results
+Our Hopper results confirm our conclusions in the TLA setting.
+All algorithms but DPO come close to matching the performance of the expert agent (2100), with CPO being the best.
+We can see SimPO is the only algorithm that significantly outperforms the bad agent (performance of 900) in the Mixed Quality setting (the $\gamma$ for SimPO was set very high, $\gamma=10$, as a lower value significantly limited performance).
+Appendix I MuJoCo Hyper-parameters
+We give the hyper-parameters we use for training.
 Parameter Value
@@
 Table 6: Hyper-parameter settings of OpenAI-ES.
+Appendix J LLM Hyper-parameters
+Parameter Value
+Number of epochs 3
+Gradient Accumulation Step 32
+Batch Size 2
+Total Batch Size 64
+Learning rate 5e-5
+LoRA Yes
+LoRA Rank 128
+LoRA Alpha 256
+Lora Dropout 0.05
+Max length 2048
+Table 7: Hyper-parameter settings for LLM Training.

### v3

@@
 Evaluating preference optimization (PO) algorithms on LLM alignment is a challenging task that presents prohibitive costs, noise, and several variables like model size and hyper-parameters.
-In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on much simpler benchmarks.
+In this work, we show that it is possible to gain insights on the efficacy of PO algorithm on simpler benchmarks.
 We design a diagnostic suite of MuJoCo tasks and datasets, which we use to systematically evaluate PO algorithms, establishing a more controlled and cheaper benchmark.
@@
 We demonstrate that our discovered PO algorithms outperform all known algorithms in the targeted MuJoCo settings.
-Finally, based on the insights gained from our MuJoCo experiments, we design a novel PO algorithm that significantly outperforms existing baselines in an LLM alignment task.
+Finally, based on the insights gained from our MuJoCo experiments, we design a PO algorithm that significantly outperform existing baselines in an LLM alignment task.
+11footnotetext: Equal contribution, order decided by coin flip.22footnotetext: Now at Amazon.
 1 Introduction
@@
 Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one.
-We use this insight to design a new PO function, Constrained Linear Policy Optimization (CoLPO) that demonstrates promising results in an LLM alingment task.
+We use this insight to design a new PO algorithm, Temporally-Aware Mirror Preference Optimization (TA-MPO), which demonstrate promising results in an LLM alignment task.
 We summarize our contributions below.
   1. 1.
-    We perform a systematic evaluation of eight of the most popular PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy.
-We see that most existing algorithms struggle when dealing with noise and mixed-quality date.
+    We perform a systematic evaluation of eight existing PO algorithms on automatically generated preference datasets with varying levels of data quality, noise levels and initial policy.
+We see that most existing algorithms struggle when dealing with noise and mixed-quality data.
   2. 2.
@@
     We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks.
-In particular, we show that our CoLPO algorithm significantly improves upon the ORPO, CPO, and SLiC-HF baselines.
-Overall, we show that it is possible to gain relevant insights for the design of LLM alignment algorithms through a systematic analysis in simple, computationally inexpensive environments such as MuJoCo.
+In particular, we show that our TA-MPO algorithm significantly improves upon the baselines.
 2 Preliminaries
+Let $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,T,\mu)$ denote an episodic Markov Decision Process, where $\mathcal{S}$ and $\mathcal{A}$ are respectively the state and action spaces, $P(s^{\prime}\mid s,a)$ is the transition probability from state $s$ to $s^{\prime}$ when taking action $a$, $r(s,a)\in[0,1]$ is the reward function, $T$ is the maximum episode length, and $\mu$ is a starting state distribution.
 A policy $\pi\in(\Delta(\mathcal{A}))^{\mathcal{S}}$, where $\Delta(\mathcal{A})$ is the probability simplex over $\mathcal{A}$, represents the behavior of an agent on an MDP, whereby at state $s\in\mathcal{S}$ the agents takes actions according to the probability distribution $\pi(\cdot\mid s)$.
@@
 Our objective is to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward of an episode, that is
-$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)$ (1)
-$:=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
+$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\tau\sim(\mu,\pi,P)}r(\tau)=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mu,a_{t},s_{t+1}}\sum_{t=0}^{T-1}r(s_{t},a_{t}),$ (1)
 where $a_{t}\sim\pi(\cdot|s_{t})$ and $s_{t+1}\sim P(\cdot|s_{t},a_{t})$.
@@
 Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
-$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w})/\eta)}{\exp(r(\tau_{w})/\eta)+\exp(r(\tau_{l})/\eta)}$ (2)
-$=\sigma((r(\tau_{w})-r(\tau_{l}))/\eta),$
-where $\sigma$ is the sigmoid function and $\eta$ is a temperature parameter.
+$\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w}))}{\exp(r(\tau_{w}))+\exp(r(\tau_{l}))}=\sigma(r(\tau_{w})-r(\tau_{l})),$ (2)
+where $\sigma$ is the sigmoid function.
 In this work, we consider an offline training setting, where the agent aim to solve the optimization problem in (1) but only has access to the the dataset $\mathcal{D}$ and cannot collect further data.
@@
 SFT is an initial alignment phase, where the policy $\pi_{0}$ is trained to imitate high-quality demonstration data.
-In particular, the starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$.
+The starting policy $\pi_{0}$ is updated to minimize the cross-entropy loss $\ell(\pi,(s_{0},\tau_{w},\tau_{l}))=-\log(\pi(\tau_{w}))$.
 We call reference policy $\pi_{\text{ref}}$ the policy obtained at the end of this procedure.
@@
 The maximum likelihood estimation problem is the one to find an estimate of the true reward function that governs how the preferences are expressed, that is
-$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}g\circ\sigma((r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l}))/\eta),$ (3)
-for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$, for an increasing function $g:\mathbb{R}\leftrightarrow\mathbb{R}$, and for $\eta\geq 0$.
-Here, typically, $g$ is set to be the logarithm function.
-The policy optimization problem is the one to maximize the expected reward, that is
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (4)
-$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$
+$\widehat{r}\in\mathop{\mathrm{argmax}}_{r_{\theta}}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\sigma(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})),$ (3)
+for a parametrized reward class $\{r_{\theta}:\theta\in\Theta\}$. The policy optimization problem is the one to maximize the expected reward, that is
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\widehat{r}(s_{t},a)-\beta D_{\mathtt{KL}}(\pi(\cdot|\tau),\pi_{\text{ref}}(\cdot|\tau))\right],$ (4)
 where $D_{\mathtt{KL}}$ represents the $\mathtt{KL}$-divergence and is introduced to prevent the policy from moving too far away from the dataset distribution.
 DPO merges these two problems by using the agent itself to implicitly represent the reward model. It consists in optimizing the objective
-$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$ (5)
-$\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$
-which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3), with $g=\log$.
-Refer to Appendix C for details.
+$\pi^{\star}$ $\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\left(\log\frac{\pi(\tau_{w})}{\pi_{\mathrm{ref}}(\tau_{w})}-\log\frac{\pi(\tau_{l})}{\pi_{\mathrm{ref}}(\tau_{l})}\right)\right)\right],$ (5)
+which is obtained by plugging the theoretical solution of (4) in the maximum likelihood problem in (3).
+Refer to Appendix˜C for details.
 Thanks to its simplicity, DPO has been widely adopted to fine-tune LLMs (Yuan et al., 2024; Jiang et al., 2024).
@@
 ORPO further simplifies the training pipeline and addresses the distribution shift issue present in DPO. It merges the SFT and DPO steps into one, optimizing the unified objective
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg{[}\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}$ (6)
-$+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg{]}$
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\bigg[\underbrace{\log\pi(\tau_{w})}_{\text{SFT}}+\lambda\underbrace{\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)}_{\text{preference optimization}}\bigg]$ (6)
 where $\operatorname{odds}_{\pi}(\tau)=\pi(\tau)/(1-\pi(\tau))$.
@@
 Additionally, the SFT term prevents pushing probability mass away from the preference dataset, addressing the distribution shift issue present in DPO.
-Research on preference optimization has been very active and many methods have been proposed. We present a summary of the most popular algorithms in Table 1 and a brief discussion in Appendix A.
+Research on preference optimization has been very active and many methods have been proposed.
+We present a summary of some among the most popular algorithms in Table˜4 and a brief discussion in Appendix A.
 Beyond these implicit algorithms, there are several other methods that explicitly solve the maximum likelihood problem in (3) and use the learned reward model to optimize the objective in (4) with an RL algorithm.
 Overall, RLHF is a superior approach and the industry standard, but is more computationally expensive and complex to implement.
-For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix F.
-Method Objective
-RRHF (Yuan et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max\left(0,-\frac{1}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)+\frac{1}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)\right)$
-SLiC-HF (Zhao et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max(0,\delta-\log\pi_{\theta}(\tau_{w}|x)+\log\pi_{\theta}(\tau_{l}|x))$
-DPO (Rafailov et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}\right)$
-IPO (Azar et al., 2024) $-\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
-CPO (Xu et al., 2024a) $\log\pi_{\theta}(\tau_{w}|x)+\log\sigma(\beta\log\pi_{\theta}(\tau_{w}|x)-\beta\log\pi_{\theta}(\tau_{l}|x))$
-ORPO (Hong et al., 2024) $\log\pi(\tau_{w})+\lambda\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)$
-R-DPO (Park et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}+(\alpha|\tau_{w}|-\alpha|\tau_{l}|)\right)$
-SimPO (Meng et al., 2024) $\log\sigma\left(\frac{\beta}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)-\frac{\beta}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)-\gamma\right)$
-Table 1: A summary of popular PO algorithms. The objective in each row is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
+For a detailed discussion and comparison between DPO-like methods and PPO, refer to Appendix H.
 2.2 Mirror Maps
 We review the concept of mirror map, which will be needed when describing our methodology.
-For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function [^1] A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$.
+For a convex set $\mathcal{X}\subseteq\mathbb{R}^{|\mathcal{A}|}$, a mirror map $h:\mathcal{X}\rightarrow\mathbb{R}$ is defined as a strictly convex, continuously differentiable and essentially smooth function***A function $h$ is essentially smooth if $\lim_{x\rightarrow\partial\mathcal{X}}\lVert\nabla h(x)\rVert_{2}=+\infty$, where $\partial\mathcal{X}$ denotes the boundary of $\mathcal{X}$.
 function that satisfies $\nabla h(\mathcal{X})=\mathbb{R}^{|\mathcal{A}|}$.
@@
 The specific class of mirror maps that we are going to use is the $\omega$-potential mirror map class, to which most mirror maps considered in the literature belong.
-Definition 2.1 ($\omega$-potential mirror map (Krichene et al., 2015)).
+Definition 2.1 ($\omega$-potential mirror map Krichene et al. (2015)).
 For $u\in(-\infty,+\infty]$, $\omega\leq 0$, an $\omega$-potential is defined as an increasing $C^{1}$-diffeomorphism $\phi:(-\infty,u)\rightarrow(\omega,+\infty)$ such that
-[…]
-For any $\omega$-potential $\phi$, we define the associated mirror map $h_{\phi}$ as
-$h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
-When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix D).
+$\lim_{x\rightarrow-\infty}\phi(x)=\omega,\;\lim_{x\rightarrow u}\phi(x)=+\infty,\;\int_{0}^{1}\phi^{-1}(x)dx\leq\infty.$
+For any $\omega$-potential $\phi$, the associated mirror map is $h_{\phi}(\pi(\cdot|s))=\sum\nolimits_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx.$
+When $\phi(x)=e^{x-1}$ we recover the negative entropy mirror map, while we recover the $\ell_{2}$-norm when $\phi(x)=2x$ (refer to Appendix˜F).
 Mirror maps in this class are simple to implement in practice, where $\mathcal{A}$ is often large, as they can be parametrized by a scalar function instead of a multi-dimentional one.
@@
 An $\omega$-potential mirror map $h_{\phi}$ induces a Bregman divergence (Bregman, 1967), which is defined as
-$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))$
-$-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
+$\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))$ $:=h_{\phi}(\pi(\cdot|s))-h_{\phi}(\pi^{\prime}(\cdot|s))-\langle\nabla h_{\phi}(\pi^{\prime}(\cdot|s)),\pi(\cdot|s)-\pi^{\prime}(\cdot|s)\rangle,$
 where $\mathcal{D}_{h_{\phi}}(\pi(\cdot|s),\pi^{\prime}(\cdot|s))\geq 0$ for all $x,y\in\mathcal{Y}$.
-When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix D).
+When $\phi(x)=e^{x-1}$, $\mathcal{D}_{h_{\phi}}$ is equivalent to the KL-divergence, while we recover the Euclidean distance when $\phi(x)=2x$ (refer to Appendix˜F).
 When the Bregman divergence is employed as a regularization term in optimization problems, tuning the mirror map allows us to control the geometry of the updates of the parameters to be optimized, determining when to take large or small updates based on the current value of the parameters.
@@
 We start by replacing the $\mathtt{KL}$-divergence penalty term in the objective in (4) with a Bregman divergence and aim to solve the problem
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}$ (7)
-$\left[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\right],$
+$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{s_{0}\sim\mathcal{D},\tau\sim(\pi,P)}\bigg[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}r(s_{t},a)-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))\bigg],$ (7)
 where $D_{h}$ is the Bregman divergence induced by a mirror map $h$.
@@
 Theorem 3.1.
-Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
-$r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (8)
+Let $h_{\phi}$ be a 0-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
+$r(\tau)=\beta\phi^{-1}(\pi^{\star}(\tau))-\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (8)
 where $c(s_{0})$ is a normalization constant that depends only on $s_{0}$.
-We provide a proof for Theorem 3.1 in Appendix C. By plugging (11) in the maximum likelihood problem in (3), we obtain the objective:
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{\mathcal{D}}\big{[}g\circ\sigma\big{(}\beta(\phi^{-1}(\pi(\tau_{w}))$ (9)
-$-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big{)}\big{]},$
+We provide a proof for Theorem˜3.1 in Appendix˜C.
+The next step is to model the reward using a classification problem based on the reward difference rather than the maximum likelihood problem in (3), as suggested by Tang et al.
+(2024).
+That is, our aim is to solve the optimization problem
+$\widehat{r}\in\mathop{\mathrm{argmax}}\nolimits_{r_{\theta}}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}g(r_{\theta}(\tau_{w})-r_{\theta}(\tau_{l})),$ (9)
+where $g$ is an increasing function. We give further details on this interpretation of reward modeling in Appendix˜D. By plugging (8) in the optimization problem in (9), we obtain the objective:
+$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{\mathcal{D}}\big[g\big(\beta(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))\big)\big],$ (10)
 where $\mathbb{E}_{\mathcal{D}}$ is equivalent to $\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}$.
-Two-step MPO
+Two-step MPO (2S-MPO)
 We can use the objective in (10) to define a class of two-step PO algorithms, which consist of a preliminary SFT phase to obtain the reference policy $\pi_{\mathrm{ref}}$ and a PO phase which optimizes (10).
-When $\phi=e^{x}$, the (9) is equivalent to (5) and we recover the DPO algorithm.
-One-step MPO
+When $\phi=e^{x}$, the (10) is equivalent to (5) and we recover DPO.
+One-step MPO (1S-MPO)
 By adding an SFT term to (10) and by setting the $\pi_{\mathrm{ref}}$ be the uniform distribution, we obtain a class of one-step PO algorithms.
 These algorithms consists in a single phase, where we optimize the objective
-$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\psi(\pi(\tau_{w}))$ (10)
-$+\lambda g\circ\sigma\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big{]},$
+$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\psi(\pi(\tau_{w}))+\lambda g\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big],$ (11)
 where $\psi$ is an $\omega$-potential.
-Here, the terms $-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ and $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))$ have canceled out due to $\pi_{\mathrm{ref}}$ being uniform.
+The term $\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))$ has canceled out due to $\pi_{\mathrm{ref}}$ being uniform.
 We note that setting $\pi_{\mathrm{ref}}$ to be the uniform distribution is equivalent to replacing the Bregman divergence penalty in (7) with the mirror map $h(\pi(\cdot|\tau))$, which enforces a form of entropy regularization.
 When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (11) recovers the ORPO objective in (6).
-The objectives in (9) and (10) allow us to implement a wide variety of preference optimization algorithms, while benefiting from a theoretical justification.
-In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform current baselines.
+Temporally-Aware MPO (TA-MPO)
+Lastly we design a variation of MPO that gradually switches from SFT to PO. TA-MPO consists of single-phase algorithms that optimize the objective
+$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[(1-\alpha(t))\psi(\pi(\tau_{w}))+\alpha(t)g\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big],$ (12)
+where $\alpha:[0,1]\rightarrow[0,1]$ is an increasing function of the percentage of training progress.
+The objectives in (10), (11), and (12) allow us to implement a variety of preference optimization algorithms, while benefiting from a theoretical justification.
+In the following, we will show that it is possible to parametrize and optimize $g$, $\psi$, and $\phi^{-1}$ to obtain new algorithms that outperform baselines.
+In Appendix˜E, we discuss why we chose Bregman divergences rather than $f$-divergences.
 3.1 Meta Learning PO objectives
 To search the space of PO algorithms we have defined, we employ a neural network parametrization for $g$, $\psi$, and $\phi^{-1}$, which we optimize using evolutionary strategies (Salimans et al., 2017).
-Similarly to (Alfano et al., 2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
-$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),$
-$e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
+Similarly to Alfano et al.
+(2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
+$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),\;e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
 where $(x)_{+}=\max(x,0)$ and $\mathop{\mathrm{clip}}(x)=\max(\min(x,1),0)$.
@@
 $\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent.
-Refer to Appendix E for further discussion and details on the ES methodology.
-In our experiments, we consider both the case where we fix $g=\log$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions.
-We also perform the evolution on both the two-step and the one-step MPO classes.
-Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
-RRHF 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
-SLiC-HF 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
-DPO 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
-IPO 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
-CPO 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
-ORPO 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
-R-DPO 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
-SimPO 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
-SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
-LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
-With $g=\log$
-1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
-1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
-2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
-2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
-With parametrized $g$
-1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
-2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
-1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
-2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
-1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
-Table 2: Three Legged Ant (TLA) Adaptation.
-Performance of popular and discovered MPO algorithms on TLA, for various dataset settings.
-For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents.
-For discovered MPO algorithms, we report performance of one-step MPO (1S-MPO) and two-step MPO (2S-MPO).
-For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters).
-We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
+Refer to Appendix˜G for further discussion and details on the ES methodology.
+We consider both the case where we fix $g=\log\sigma$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions.
+We perform the evolution on both the two-step and one-step MPO classes.
+Figure 1: Data generation for MuJoCo experiments.
+For each pair of trajectories, which share the same starting state, two agents are chosen based on the data generating strategy.
+The trajectories are then judged by a Bradley-Terry Judge and, in the noisy setting, the labels are flipped with probability $\varepsilon$.
 4 MuJoCo Experiments
 Our first set of experiments is carried out on continuous RL tasks in MuJoCo.
-In particular, we show the performance of all the algorithms presented in Table 1 across several settings and we compare it with the performance of our discovered objectives.
+In particular, we show the performance of all the algorithms presented in Table˜4 across several settings and we compare it with the performance of our discovered objectives.
 To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries.
-We report our hyper-parameters in Appendix I.
+We provide an implementation of our methodology here and report hyper-parameters in Appendix˜K.
 4.1 Tasks
@@
 A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward.
-We also consider training a randomly initialized agent on the Hopper environment, using a preference dataset with 5120 pairs of trajectories.
-This setting addresses the case where a new behavior has to be learned completely from the preference dataset and no prior knowledge of the task is available.
-Due to space constraints, we report the results of the experiments for this task in Appendix H.
 We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
   • Base dataset: for each pair of trajectories, one is generated by the original agent and one by the target agent.
-  • Noisy dataset: same as the base dataset but each label given by the judge is flipped with probability $\varepsilon$.
+  • Noisy dataset: same as the base dataset but each chosen/rejected pair of labels given by the judge is flipped with probability $\varepsilon$.
   • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one.
-The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between a trajectory of the target agent and one of the original agent, and $25\%$ comparisons between two trajectories of the original agent.
-Figure 1: SimPO
-Figure 2: Discovered 1S-MPO (shuffled)
-Figure 3: ORPO
-Figure 4: Discovered 1S-MPO (noisy)
-Figure 5: Plots of the absolute value of the gradient of SimPO, ORPO, and of two discovered objectives, i.e.
-1S-MPO (shuffled) and 1S-MPO (noisy, $\epsilon=0.3$).
-The dots are sampled datapoints from the training distribution, with lighter points representing more recent datapoints during the course of training.
-In the bottom right and corner of each plot, SimPO’s and ORPO’s gradients go to zero, while the gradient of our discovered loss functions keep increasing.
-The effect of this on training is evident observing the datapoints: in ORPO and SimPO, the likelihood of the rejected points is not minimized, leading to decreased performance.
+The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between trajectories of different agents, and $25\%$ comparisons between two trajectories of the original agent.
+We also consider training a randomly initialized agent on the Hopper environment.
+This setting addresses the case where a behavior has to be learned from the preference dataset and there is no prior knowledge of the task available.
+We report the results of the experiments for this task in Table˜6.
 4.2 Results
-We provide the results of our experiments for TLA in  Table 2, which reports the performance of several PO algorithm and of our discovered objectives.
+We provide the results of our experiments for TLA in  Table˜1, which reports the performance of several PO algorithm and of our discovered objectives.
 We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters.
-All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
-First of all, we notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a large drop in performance in the mixed-quality and noisy settings.
+All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO, SimPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
+We provide an additional noisy setting ($\varepsilon=0.2$) and performance for other existing algorithms in Table˜5 in Appendix I.1.
+We notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a drop in performance in mixed-quality and noisy settings.
 Importance of SFT
-The first group within Table 2 shows that SFT plays a key role in the performance across different settings.
+The first group within Table˜1 shows that SFT plays a key role in the performance across different settings.
 SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings.
@@
 CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
+Table 1: Three Legged Ant (TLA).
+Performance of existing and discovered MPO algorithms on TLA, for various dataset settings.
+For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents.
+For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters).
+We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
+Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.3$)
+RRHF (Yuan et al., 2023) 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 330 $\pm$ 552
+SLiC-HF (Zhao et al., 2023) 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1135 $\pm$ 224
+DPO (Rafailov et al., 2024) 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 1519 $\pm$140
+IPO (Azar et al., 2024) 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 1133 $\pm$ 115
+CPO (Xu et al., 2024a) 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2000 $\pm$35
+ORPO (Hong et al., 2024) 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 1953 $\pm$ 37
+R-DPO (Park et al., 2024) 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1667 $\pm$ 24
+SimPO (Meng et al., 2024) 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 -3828 $\pm$ 341
+SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2049 $\pm$33
+Our algorithms $\downarrow$
+LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 1569 $\pm$ 156
+With $g=\log\sigma$
+1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -3967 $\pm$ 382
+1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3279 $\pm$ 83
+2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 1687 $\pm$ 58
+2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 1569 $\pm$ 122
+With parametrized $g$
+1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 2681 $\pm$ 251
+2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2253 $\pm$ 125
+1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 1771 $\pm$ 107
+2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2074 $\pm$ 136
+1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3417 $\pm$ 82
+Temporally-aware
+TA-MPO (1) 3577 $\pm$ 45 2730 $\pm$ 63 3106 $\pm$ 62 1971 $\pm$ 35
+TA-MPO (2) 3625 $\pm$ 49 3088 $\pm$ 69 3443 $\pm$ 53 1988 $\pm$ 39
+TA-MPO (3) 3352 $\pm$ 57 2256 $\pm$ 44 2725 $\pm$ 46 1923 $\pm$ 27
 Keep optimizing
@@
 To further stress this point, we consider the objective
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big{]},$
+$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big],$
 which we call Linear Preference Optimization (LPO).
-LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table 2, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
+LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table˜1, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
 Discovered objectives
-Table 2 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it.
+Table˜1 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it.
 We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings.
@@
 This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
-When exploring the one-step MPO class with $g=\log$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
-$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\alpha\log(\pi(\tau_{w}))$
-$+\lambda\log\sigma\left(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\right)\big{]},$
+Figure 2: SimPO
+Figure 3: ORPO
+Figure 4: Discovered 1S-MPO (shuffled)
+Figure 5: Discovered 1S-MPO (noisy)
+Figure 6: Absolute value of the gradient of SimPO, ORPO, 1S-MPO (shuffled), and 1S-MPO (noisy, $\epsilon=0.3$). The dots are sampled datapoints from the training distribution.
+When exploring the one-step MPO class with $g=\log\sigma$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
+$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\alpha\log(\pi(\tau_{w}))+\lambda\log\sigma\big(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\big)\big],$
 where the coefficients $\alpha$ and $\beta$ depend on the setting.
@@
 These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories.
-When we search the two-step MPO class with $g=\log$, we recover DPO.
-Figure 5 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$.
+When we search the two-step MPO class with $g=\log\sigma$, we recover DPO.
+Figure˜6 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$.
 In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings.
@@
 We highlight once more that both discovered objectives advocate to keep optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
-For additional results and analysis for the Hopper task, we refer the reader to Appendix H.
-WR $s_{\text{WR}}$ LCWR $s_{\text{LCWR}}$ avg length
-SLiC-HF 16.77 1.32 13.18 0.28 3016
-ORPO 24.47 1.52 19.25 0.29 1679
-CPO 25.96 1.54 21.07 0.25 1623
-CoLPO 31.94 1.65 24.01 0.21 2391
+4.3 Including temporal awareness
+We build an algorithm within the TA-MPO family using the insights gained in the previous section.
+In particular, we keep the standard SFT loss, which was rediscovered in all our experiments, and use SimPO for the PO component of TA-MPO, given its high performance and its ability to continue optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
+That is, we define the objective
+$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{\mathcal{D}}\big[(1-\alpha(t))\log\pi_{\theta}(\tau_{w}|x)+\alpha(t)\log\sigma(\beta(\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x))-\gamma)\big].$ (13)
+We try three expressions for $\alpha$, designed to put most of the weight on the SFT component of (13) at the start of training and switch to PO towards the end of training:
+$(1):\;\alpha(t)=t;\quad(2):\;\alpha(t)=t^{2};\quad(3):\;\alpha(t)=\sigma(20(x-0.75)).$
+Table˜1 shows promising results for temporally-aware PO algorithms, as the objective in (13) with the version (2) of $\alpha$ matches or surpasses all human-designed algorithms in all settings.
 Table 2: AlpacaEval LLM results.
-We report Win Rate (WR), standard error for Win Rate ($s_{\text{WR}}$), Length Controlled Win Rate (LCWR), standard error for LCWR ($s_{\text{LCWR}}$) and average length of responses.
-We trained gemma-7b on dpo-mix-7k for 3 epochs.
-We then compared our model to GPT-4 using Llama-3.3-70B-Instruct as judge.
-All algorithms use the same $\beta=0.5$ and the same learning rate of $55e-$ 5.
-In the case of CoLPO and SLiC-HF, we used the same cutoff rate at 5.
+We report win-rates and standard error (length controlled win-rates and standard error in parenthesis) for three combinations of base model and preference dataset.
+We report in bold font the highest winrate (length-controlled winrate) for each column.
+gemma-7bhttps://huggingface.co/google/gemma-7b, dpo-mix-7khttps://huggingface.co/datasets/argilla/dpo-mix-7k gemma-7b, capybara-7k§§§https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized mistral-7bhttps://huggingface.co/mistralai/Mistral-7B-v0.3, dpo-mix-7k
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 5 Experiments: LLM transfer
 We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting.
-In particular, we found that all our discovered loss functions kept optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, while all algorithms in Table 1 have a PO term whose gradient quickly reaches 0 when $\pi(\tau_{w})>\pi(\tau_{l})$.
-Using this insight, we design a new PO objective, which we call Constrained Linear Preference Optimization (CoLPO) and define as
-$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(\tau_{w}|x)$
-$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)<\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),0)$
-$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)>\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),\gamma)\big{]},$
-where $\beta\geq 0$, $\gamma\geq 0$ and $\alpha\leq 0$ are hyperparameters.
-The CoLPO objective keeps optimizing the policy when $\pi(\tau_{w})>\pi(\tau_{l})$, but only if $\log\pi_{\theta}(\tau_{w}|x)$ is over a certain threshold.
-This constraint is placed to stop the optimizer from increasing the difference $\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))$ by descreasing both probabilities.
-We evaluate CoLPO on an LLM fine-tuning task, where we fine-tune the gemma-7b (Team et al., 2024) model on the dpo-mix-7k [^2] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset for 3 epochs.
-Being a real-world preference dataset, dpo-mix-7k naturally presents characteristics of the mixed-quality and the noisy settings.
-To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the CoLPO objective.
-We evaluate the tuned LLM against GPT-4, using the AlpacaEval library  (Li et al., 2023).
-All models were trained using the default hyper-parameters, with $\beta=0.5$, $\gamma=5$ (for CoLPO and SLiC-HF), and $\alpha=-1$ (for CoLPO).
-For details on hyperparameters and training settings consult Appendix J.
-Table 3 reports the results of CoLPO, as well as the results for other one-step algorithms, and shows that CoLPO largely outperforms existing algorithms, confirming the validity of the insights gained in the MuJoCo simulations.
-Lastly, we report that the the log-probabilities for the final training batch of chosen and rejected responses are -1.17 and -11.38 respectively for CoLPO, compared to -0.61 and -2.98 for ORPO, respectively.
-This results further confirms that CoLPO is effective in decreasing the probability of rejected responses and mantaining high the probability of chosen responses.
+In particular, we test the TA-MPO algorithm in (13), which is designed to continue to optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, as all our discovered algorithms do.
+We evaluate TA-MPO on LLM alignment, comparing its performance against baselines for three combinations of base model and preference dataset, as shown in Table˜2.
+To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the TA-MPO objective in (13).
+We evaluate the tuned LLMs against GPT-4, using the AlpacaEval library  (Li et al., 2023) and Llama-3.1-70B-Instruct as a judge.
+For all combinations of starting LLM, dataset, and PO algorithm, we perform 4 update epochs and set the learning rate to 5e-5 and $\beta$ to 0.05.
+In the case of DPO and SimPO, we performed 3 epochs of SFT, with learning rate 5e-5, and 1 epoch of DPO/SimPO, with learning rate 5e-7 and $\beta=0.05$.
+Refer to Appendix˜L for further details.
+Table˜2 shows the effectiveness of the TA-MPO objective in (13), which presents a high winrate for all schedules of $\alpha$.
+In particular, TA-MPO with the sigmoid schedule for $\alpha$ has the highest winrate in all tasks and the highest length controlled winrate in two out of three tasks.
+Additionally, TA-MPO requires only one stage of training, while DPO and SimPO require two, i.e.
+SFT and PO.
+We also tested the static algorithms discovered in the MuJoCo environment.
+Those with $g=\log\sigma$, which rediscovered CPO, exhibited performance similar to CPO itself.
+In contrast, algorithms with a parameterized $g$ function, which learned to greedily optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, performed poorly.
+We hypothesize that these algorithms overfit to the TLA task, where our datasets sufficiently cover the state-action space.
+On the other hand, in LLM tuning—where data is sparse relative to the environment—greedy methods are more susceptible to over-optimization.
+At the same time, the objectives discovered on TLA are only used to logits within -14 and -8, as shown in Figure˜6, and have less regular shape outside of this subset.
+In Table˜3, we report additional results that allow us to better understand the influence of different components of TA-MPO.
+In particular, we have tested SimPO with $\gamma=0$, SimPO with $\gamma>0$, and SFT + SimPO with $\gamma>0$, where the last one consists of a one-step objective made of the addition between the SFT and the SimPO loss.
+Table˜3 shows that having a large $\gamma$, which encourages large updates even when $\pi(\tau_{w})>\pi(\tau_{l})$, or adding an SFT term to the PO loss are both helpful in improving performance.
+The performance of TA-MPO reported in Table˜2 is still the highest, which means that temporal awareness also contributes to a better performance.
+We can therefore confirm that each of the insights we gained in the MuJoCo benchmark transfers to the LLM alignment setting.
+Table 3: We report win-rates and standard error on AlpacaEval (length controlled win-rates and standard error in parenthesis) for gemma-7b and two preference dataset.
+Algorithm $\gamma$ dpo-mix-7k capybara-7k
+SimPO 0 30.50 (26.06) 27.52 (25.12)
+SimPO 1 30.62 (25.35) 27.52 (25.54)
+SimPO 2 32.50 (25.80) 27.95 (25.71)
+SimPO 5 33.42 (28.22) 27.02 (25.07)
+SimPO 10 33.60 (27.12) 27.39 (25.00)
+SFT+SimPO 1 30.31 (26.55) 35.07 (28.25)
 6 Conclusion
@@
 Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$.
-To address this, we proposed a simple objective, CoLPO, that avoids such loss truncation while preventing the optimizer from decreasing the probability of chosen datapoints.
+Using this insight, we proposed a temporally-aware algorithm, TA-MPO, that avoids such loss truncation and gradually switches from the SFT step to the PO step.
 We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
-Appendix A Popular PO Algorithms
+Acknowledgments and Disclosure of Funding
+Carlo Alfano and Silvia Sapora are supported by the Engineering and Physical Sciences Research Council EP/W524311/.
+Jakob Foerster is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC).
+Jakob Foerster is also supported by the JPMC Research Award and the Amazon Research Award.
+Patrick Rebeschini is funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number EP/Y028333/1].
+Yee Whye Teh acknowledges support from the Ministry of Digital Development and Information (MDDI) under the Singapore Global AI Visiting Professorship Program (Award No.
+AIVP-2024-002).
+Appendix A Baseline PO Algorithms
+Table 4: Overview of popular PO algorithms. The objective is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
+Method Objective
+RRHF (Yuan et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max\left(0,-\frac{1}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)+\frac{1}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)\right)$
+SLiC-HF (Zhao et al., 2023) $\lambda\log\pi_{\theta}(\tau_{w}|x)-\max(0,\delta-\log\pi_{\theta}(\tau_{w}|x)+\log\pi_{\theta}(\tau_{l}|x))$
+DPO (Rafailov et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}\right)$
+IPO (Azar et al., 2024) $\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
+CPO (Xu et al., 2024a) $\log\pi_{\theta}(\tau_{w}|x)+\log\sigma(\beta\log\pi_{\theta}(\tau_{w}|x)-\beta\log\pi_{\theta}(\tau_{l}|x))$
+ORPO (Hong et al., 2024) $\log\pi(\tau_{w})+\lambda\log\sigma\left(\log\left(\operatorname{odds}_{\pi}(\tau_{w})\right)-\log\left(\operatorname{odds}_{\pi}(\tau_{l})\right)\right)$
+R-DPO (Park et al., 2024) $\log\sigma\left(\beta\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\beta\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}+(\alpha|\tau_{w}|-\alpha|\tau_{l}|)\right)$
+SimPO (Meng et al., 2024) $\log\sigma\left(\frac{\beta}{|\tau_{w}|}\log\pi_{\theta}(\tau_{w}|x)-\frac{\beta}{|\tau_{l}|}\log\pi_{\theta}(\tau_{l}|x)-\gamma\right)$
 These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT.
@@
 Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024).
-An approach particularly relevant to our method is DiscoPOP by (Lu et al., 2024), which leverages an LLM to discover objective functions for LLM tuning.
-They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by (Tang et al., 2024).
+An approach particularly relevant to our method is DiscoPOP by Lu et al.
+(2024), which leverages an LLM to discover objective functions for LLM tuning.
+They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by Tang et al.
+(2024).
 Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space.
@@
 Generalisations of DPO
-A generalization of DPO alternative to ours is $f$-DPO, developed by (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function.
+A generalization of DPO alternative to ours is $f$-DPO (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function.
 We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa.
 They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms.
-(Huang et al., 2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
+Huang et al.
+(2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
 Appendix C Proof of Theorem 3.1
-We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
-Theorem C.1 (Theorem 3.1).
-Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
+We provide here a proof for our main result, i.e. Theorem˜3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
+Theorem C.1 (Theorem˜3.1).
+Let $h_{\phi}$ be a 0-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
 $r(\tau)=\phi^{-1}(\pi^{\star}(\tau))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau))+c(s_{0}),$ (14)
@@
 We use the stationarity condition to obtain the equation
-$\nabla_{\pi(\tau)}$ $\Bigg{[}\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))$
-$\quad-\lambda\left(\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\pi(\tau^{\prime})-1\right)+\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg{]}$
-$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
+$\nabla_{\pi(\tau)}$ $\Bigg[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))-\lambda\hskip-8.5359pt\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\hskip-8.5359pt\pi(\tau^{\prime})-\lambda+\hskip-8.5359pt\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\hskip-8.5359pt\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg]$
+$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
 for all initial states $s_{0}\in\mathcal{S}$ and for all trajectories $\tau$ starting from $s_{0}$. Rearranging, we obtain that
-$\pi(\tau)=\phi\big{(}r(\tau)+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)\big{)}.$
-Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 00-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$.
+$\pi(\tau)=\phi\big((r(\tau)+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau))/\beta\big).$
+Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 0-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$.
 Invoking the complementary slackness condition, whereby $\alpha(\tau)\pi(\tau)=0$ for all trajectories $\tau$, we have that $\alpha(\tau)=0$ for all trajectories $\tau$.
 Therefore, we have that
-$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
+$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
 The theorem statement is obtained by rearranging the last equation and denoting $c(s_{0})=\lambda$
-Appendix D Further discussion of $\omega$-potentials
+Appendix D Reward Modeling
+In Equation (9), we utilize the interpretation of reward modeling as a binary classification problem given by Tang et al.
+(2024), which we summarize here.
+Let $z=(\tau_{1},\tau_{2})$ be a pair of trajectories and $\ell\in\{-1,1\}$ be the associated label that states whether $\tau_{1}$ is preferred to $\tau_{2}$ ($\ell=1$) or not ($\ell=-1$).
+We want to find a function $\hat{\ell}(z)\in\mathbb{R}$ such that $\mathop{\mathrm{sign}}(\hat{\ell}(z))$ is a good estimate of $\ell$.
+For a dataset $\mathcal{D}=\{z_{i},\ell_{i}\}_{i=1}^{N}$, the classification loss (or 0-1 loss) is
+$L(\widehat{\ell},\mathcal{D})=\mathbb{E}_{\mathcal{D}}\left[1-\text{sign}\left(\hat{\ell}(z)\cdot\ell\right)\right],$ (15)
+which is often approximated with a surrogate
+$L_{f}(\widehat{\ell},\mathcal{D})=\mathbb{E}_{\mathcal{D}}\left[f\left(\hat{\ell}(z)\cdot\ell\right)\right],$ (16)
+for a function $f:\mathbb{R}\to\mathbb{R}$.
+This approximation is possible because, when $f$ is decreasing (or convex), (15) and (16) have the same minimizer, as we prove in the following.
+Denote $p_{1}(z)=\mathbb{P}(\ell=1|z)$, then the conditional surrogate loss at $z$ is
+$L_{f}(\widehat{\ell},x)=p_{1}(z)f\left(\hat{\ell}(z)\right)+(1-p_{1}(z))f\left(-\hat{\ell}(z)\right).$ (17)
+The minimizer $\widehat{\ell}^{\star}$ of (16) is such that $\widehat{\ell}^{\star}(z)$ minimizes (17).
+While the minimizer to (17) might not be computable explicitly, we can show $\widehat{\ell}^{\star}(z)>0\iff p_{1}(z)>1/2$ when $f$ is a decreasing function, meaning that $\widehat{\ell}^{\star}$ is also the minimizer of (15).
+Firstly, we have that
+$L_{f}(\widehat{\ell},x)-L_{f}(-\widehat{\ell},x)$ $=p_{1}(z)f\left(\hat{\ell}(z)\right)+(1-p_{1}(z))f\left(-\hat{\ell}(z)\right)$ (18)
+$\quad-p_{1}(z)f\left(-\hat{\ell}(z)\right)-(1-p_{1}(z))f\left(\hat{\ell}(z)\right)$
+$=\left(2p_{1}(z)-1\right)\left(f\left(\hat{\ell}(z)\right)-f\left(-\hat{\ell}(z)\right)\right)$
+Since $f$ is decreasing, we have for all $\widehat{\ell}(z)>0$ that $f(\widehat{\ell}(z))<f(-\widehat{\ell}(z))$.
+Plugging this into (18), we obtain that, if $p_{1}(z)>1/2$, $L_{f}(\widehat{\ell},x)<L_{f}(-\widehat{\ell},x)$ for all $\widehat{\ell}(z)>0$.
+Therefore, the minimizer $\widehat{\ell}^{\star}(z)$ of (17) must be positive.
+The opposite can be proved in the same manner.
+Equation (9) can be obtained by setting f=gf=-g, for an increasing function $g$, and
+$\widehat{\ell}(z)=\widehat{r}(\tau_{1})-\widehat{r}(\tau_{2}).$
+Appendix E Bregman divergences vs $f$-divergences
+In this section, we discuss why we use Bregman divergences rather than $f$-divergences in (7).
+Firstly, we note that, if the reference policy $\pi_{\mathrm{ref}}$ is set to be the uniform distribution, Bregman divergences and $f$-divergences generate equivalent families of algorithms.
+Let $\phi$ be a 0-potential.
+Using a reasoning similar to that of Theorem˜3.1, we have the following cases:
+Bregman divergence $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l})),$
+${\scriptsize\text{+uniform }\pi_{\mathrm{ref}}}$ $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l})),$
+$f$-divergence $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w})/\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l})/\pi_{\mathrm{ref}}(\tau_{l})),$
+${\scriptsize\text{+uniform }\pi_{\mathrm{ref}}}$ $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(|\mathcal{T}|(\pi^{*}(\tau_{w})))-\phi^{-1}(|\mathcal{T}|(\pi^{*}(\tau_{l})))$
+$\qquad\qquad\qquad\;\;=\tilde{\phi}^{-1}(\pi^{*}(\tau_{w}))-\tilde{\phi}^{-1}(\pi^{*}(\tau_{l})),$
+where $\mathcal{T}$ is the set of all trajectories and $\tilde{\phi}^{-1}(x)=\phi^{-1}(|\mathcal{T}|x)$.
+The two resulting expressions are equivalent, meaning that our 1S-MPO class also includes the case of $f$-divergences.
+Our experiment on the 1S-MPO class therefore explore both Bregman divergences and $f$-divergences with ES and Table˜1 reports the performance of the best divergence found within both divergence classes.
+If the reference policy is not uniform, then the two algorithmic classes have the KL divergence as the only intersection.
+In this setting, we have chosen to focus on Bregman divergences as Wang et al.
+(2023) have already shown that, among the $f$-divergences they considered, the KL-divergence typically offers superior alignment performance.
+Since we wanted to explore a class of algorithms with the objective of finding better alternatives to the KL-divergence, we decided to consider a different generalization of the KL-divergence, i.e.
+Bregman divergences.
+In our experiments with $g=\log$, we discovered that the KL-divergence is optimal also within Bregman divergences.
+On the other hand, we found out that significant improvements in performance come from modifying $g$ rather than the KL-divergence.
+We have observed similar results in preliminary MuJoCo experiments where we replaced the Bregman divergence with an $f$-divergence, that is that the KL-divergence is optimal among $f$-divergences.
+Appendix F Further discussion of $\omega$-potentials
 We show here two examples of Bregman divergence induced by an $\omega$-potential mirror map, that is when $\phi(x)=e^{x-1}$ and when $\phi(x)=x$.
 If $\phi(x)=e^{x-1}$, the associated mirror map is defined as
-[…]
+$h_{\phi}(\pi(\cdot|s))$ $=\sum_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx=\sum_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}(\log(x)+1)dx$
 $=\sum_{a\in\mathcal{A}}\pi(a\mid s)\log(\pi(a\mid s))-\pi(a\mid s)+\pi(a\mid s)$
@@
 which is the definition of the $\mathtt{KL}$-divergence. If $\phi(x)=2x$, the associated mirror map is defined as
-[…]
-$=\sum_{a\in\mathcal{A}}\pi(a\mid s)^{2},$
+$h_{\phi}(\pi(\cdot|s))$ $=\sum_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}\phi^{-1}(x)dx=\sum_{a\in\mathcal{A}}\int_{1}^{\pi(a\mid s)}2xdx=\sum_{a\in\mathcal{A}}\pi(a\mid s)^{2},\$
 which is the $\ell_{2}$-norm. Plugging this expression in the definition of Bregman divergence we obtain
-$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
-$=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})$
-$=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
+$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
 which is the definition of the Euclidean distance.
-Appendix E Further discussion on Evolution Strategies
+Appendix G Further discussion on Evolution Strategies
 Evolution Strategies (ES) represent a powerful, backpropagation-free method for optimizing complex functions, that has been particularly successful in the context of long-horizon, noisy, and bi-level optimization tasks such as RL and meta-RL.
@@
 This feature makes ES well-suited for tasks with long computational graphs, for instance algorithms with many updates, where, due to memory constraints, traditional gradient-based methods have to resort to gradient truncation, introducing bias (Werbos, 1990; Metz et al., 2022; Liu et al., 2022).
-In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section 3.1, so that an agent trained using the objective in (10) achieves the highest value.
+In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section˜3.1, so that an agent trained using the objective in (11) achieves the highest value.
 Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (11) when using the parametrized $\psi$ and $\phi^{-1}$.
@@
 where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
-Appendix F Further discussion on Online vs Offline Methods
+Appendix H Further discussion on Online vs Offline Methods
 In the domain of RL and preference optimization, the choice between online and offline algorithms presents a critical trade-off, influencing computational efficiency, data requirements, and generalization capabilities.
@@
 To mitigate this issue, DPO-iter (Xu et al., 2024b), which incorporates online data, has been proposed as a potential solution.
-Appendix G Further discussion on Meta-Learning Algorithms
+Appendix I MuJoCo Additional Results
+I.1 TLA
+Table 5: Three Legged Ant (TLA).
+Performance of existing and discovered MPO algorithms on TLA.
+For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents.
+For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters).
+We report in bold the highest (or two highest if their confidence interval overlaps) average performance, for each setting.
+Base Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
+RRHF 2789 $\pm$ 285 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
+SLiC-HF 3255 $\pm$ 66 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
+DPO 3528 $\pm$ 58 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
+IPO 3618 $\pm$ 44 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
+CPO 3450 $\pm$ 55 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
+ORPO 3087 $\pm$ 322 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
+R-DPO 2606 $\pm$ 65 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
+SimPO 3683 $\pm$78 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
+SFT 3287 $\pm$ 62 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
+KTO 1534 $\pm$ 34 1551 $\pm$ 31 1531 $\pm$ 49 1442 $\pm$ 35
+f-DPO (Jensen-Shannon) 3621 $\pm$ 50 3192 $\pm$ 76 2494 $\pm$ 101 1633 $\pm$ 123
+Our algorithms $\downarrow$
+LPO 3774 $\pm$ 102 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
+With $g=\log\sigma$
+1S-MPO (mixed-quality) 3206 $\pm$ 330 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
+1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
+2S-MPO (mixed-quality) 3595 $\pm$ 57 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
+2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
+With parametrized $g$
+1S-MPO (mixed-quality) 3560 $\pm$ 333 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
+2S-MPO (mixed-quality) 3736 $\pm$ 51 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
+1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
+2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
+1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
+We display additional results for the TLA task in Table˜5.
+With respect to Table˜1, we replace the mixed-quality setting with a noisy setting where the noise parameter $\varepsilon$ is set to 0.2.
+We also include the KTO (Ethayarajh et al., 2024) and $f$-DPO (Jensen-Shannon) (Wang et al., 2023) algorithms.
+I.2 Simplified expression for discovered objectives
+Below, we report a simplified version of the objectives discovered for the shuffled and noisy ($\epsilon=0.3$) settings, when $g$ is parametrized:
+$\mathcal{L}(\tau_{w},\tau_{l})=$ $0.82\mathcal{L}_{\mathrm{SFT}}(\tau_{w})+1.7(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))$
+$+0.33(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{2}+0.36(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{3}$
+$\mathcal{L}(\tau_{w},\tau_{l})=$ $0.82\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$
+$+\max(1.39(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{2},0.12(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
+I.3 Hopper Tasks
+Table 6: Hopper. Perfomance of existing MPO algorithms on the Hopper setting. The agent is randomly initialised.
+Base Mixed Quality Noisy ($\epsilon=0.1$)
+DPO 1796 $\pm$ 78 458 $\pm$ 82 693 $\pm$ 131
+IPO 2049 $\pm$ 13 1606 $\pm$ 91 739 $\pm$ 118
+CPO 2078 $\pm$ 12 1078 $\pm$ 35 1813 $\pm$ 33
+ORPO 2022 $\pm$ 15 1039 $\pm$ 20 1710 $\pm$ 33
+SimPO 2027 $\pm$ 15 1460 $\pm$ 94 1794 $\pm$ 65
+We consider a second set of simulations on MuJoCo, based on the Hopper environment.
+As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries.
+We report our hyper-parameters in Appendix˜K.
+Differently from the TLA task, we consider a setting where the agent is randomly initialized and needs to learn a policy from scratch.
+On Hopper, we train an agent with an expected cumulative reward of 2100 (the expert agent) and an agent with an expected cumulative reward of 900 (the bad agent).
+We generate the preference datasets in the same way we do for TLA, with the exceptions that the number of rows is 5120 and the trajectories are generated by either the expert or the bad agent.
+We consider the same three variations of the preference datasets used in TLA, where the expert agent corresponds to the target agent, and the bad agent corresponds to the original agent.
+We include more data compared to the TLA setting as it takes more datapoints for the agent to learn from scratch rather than to adapt to a slightly different objective (like in the TLA case).
+Our Hopper results confirm our conclusions in the TLA setting.
+All algorithms but DPO come close to matching the performance of the expert agent (2100), with CPO being the best.
+We can see SimPO is the only algorithm that significantly outperforms the bad agent (performance of 900) in the Mixed Quality setting (the $\gamma$ for SimPO was set very high, $\gamma=10$, as a lower value significantly limited performance).
+I.4 Further MuJoCo Analsys
+In addition to the noisy dataset, we also considered a bad judge setting, where the judge would be more likely to swap the label of a pair of trajectories if their ground truth rewards were closer to each other.
+This is practically implemented as an increase in the temperature of the Bradley-Terry judge.
+However, we did not notice significantly different results compared to the simple noisy setting, therefore detailed results are not reported.
+Appendix J Further discussion on Meta-Learning Algorithms
 Meta-learning, or “learning to learn”, has been extensively employed to automate the design of algorithms that can either adapt rapidly with minimal data samples or generalize effectively to unseen data, tasks, or environments.
@@
 In our setting, the inner loop is the offline preference optimization algorithm, while the outer loop is the agent evaluation on the environment (online) and the update of the meta-parameters $\zeta$.
-Base Mixed Quality Noisy $\epsilon=0.1$
-DPO 1796 $\pm$ 78 458 $\pm$ 82 693 $\pm$ 131
-IPO 2049 $\pm$ 13 1606 $\pm$ 91 739 $\pm$ 118
-CPO 2078 $\pm$ 12 1078 $\pm$ 35 1813 $\pm$ 33
-ORPO 2022 $\pm$ 15 1039 $\pm$ 20 1710 $\pm$ 33
-SimPO 2027 $\pm$ 15 1460 $\pm$ 94 1794 $\pm$ 65
-Table 4: Hopper. Perfomance of popular MPO algorithms on the Hopper setting. The agent is randomly initialised.
-Appendix H MuJoCo Additional Results
-Below, we report a simplified version of the discovered loss on the shuffled and noisy $\epsilon=0.3$ datasets:
-$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (13)
-$+1.4\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
-$+0.22\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2}$
-$+0.2\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{3}$
-$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (14)
-$+\max((1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2},$
-$0.1\cdot(1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))))$
-We consider a second set of simulations on MuJoCo, based on the Hopper environment.
-As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries.
-We report our hyper-parameters in Appendix I.
-H.1 Hopper Tasks
-Differently from the TLA task, we consider a setting where the agent is randomly initialized and needs to learn a policy from scratch.
-On Hopper, we train an agent with an expected cumulative reward of 2100 (the expert agent) and an agent with an expected cumulative reward of 900 (the bad agent).
-We generate the preference datasets in the same way we do for TLA, with the exceptions that the number of rows is 5120 and the trajectories are generated by either the expert or the bad agent.
-We consider the same three variations of the preference datasets used in TLA, where the expert agent corresponds to the target agent, and the bad agent corresponds to the original agent.
-We include more data compared to the TLA setting as it takes more datapoints for the agent to learn from scratch rather than to adapt to a slightly different objective (like in the TLA case).
-H.2 Further MuJoCo Analsys
-In addition to the noisy dataset, we also considered a bad judge setting, where the judge would be more likely to swap the label of a pair of trajectories if their ground truth rewards were closer to each other.
-This is practically implemented as an increase in the temperature of the Bradley-Terry judge.
-However, we did not notice significantly different results compared to the simple noisy setting, therefore detailed results are not reported.
-H.3 Results
-Our Hopper results confirm our conclusions in the TLA setting.
-All algorithms but DPO come close to matching the performance of the expert agent (2100), with CPO being the best.
-We can see SimPO is the only algorithm that significantly outperforms the bad agent (performance of 900) in the Mixed Quality setting (the $\gamma$ for SimPO was set very high, $\gamma=10$, as a lower value significantly limited performance).
-Appendix I MuJoCo Hyper-parameters
+Appendix K MuJoCo Hyper-parameters
 We give the hyper-parameters we use for training.
+The hyper-parameters specific to each algorithm are tuned for each task-data type combination.
+All the experiments were conducted on 4 NVIDIA L40S GPUs.
+Table 7: Hyper-parameter settings for PO.
 Parameter Value
@@
 Max gradient norm 1.3
-$\lambda$ 0.5
-Table 5: Hyper-parameter settings for PO.
+Table 8: Hyper-parameter settings of OpenAI-ES.
 Parameter Hopper TLA
@@
 Learning rate 0.02 0.02
-Table 6: Hyper-parameter settings of OpenAI-ES.
-Appendix J LLM Hyper-parameters
+Appendix L LLM Hyper-parameters
+We give the hyper-parameters we use for LLM training. All the experiments were conducted on 4 NVIDIA L40S GPUs.
+Table 9: Hyper-parameter settings for LLM Training.
 Parameter Value
-Number of epochs 3
 Gradient Accumulation Step 32
@@
 Total Batch Size 64
-Learning rate 5e-5
 LoRA Yes
@@
 Max length 2048
-Table 7: Hyper-parameter settings for LLM Training.

### v4

@@
 Learning from human preferences (Christiano et al., 2017) is a paradigm which enables the alignment of machine learning systems to relative human preferences, without requiring access to absolute rewards.
-While the framework was developed for robotic and games applications with experiments on MuJoCo simulations and Atari (Akrour et al., 2012; Biyik & Sadigh, 2018; Ibarz et al., 2018), this paradigm has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023).
+While the framework was developed for robotic and games applications with experiments on MuJoCo simulations and Atari (Akrour et al., 2012; Biyik and Sadigh, 2018; Ibarz et al., 2018), this paradigm has been successfully applied to Large Language Models (Team et al., 2023; Achiam et al., 2023).
 In particular, fine-tuning pre-trained LLMs with human preferences has become a popular strategy to adapt them to specific tasks and to improve their safety and helpfulness.
@@
 Moreover, we introduce a framework for finding PO algorithms.
-Specifically, we define a class of PO algorithms based on mirror descent (Nemirovski & Yudin, 1983), which generalizes DPO and ORPO for particular choices of the mirror map.
+Specifically, we define a class of PO algorithms based on mirror descent (Nemirovski and Yudin, 1983), which generalizes DPO and ORPO for particular choices of the mirror map.
 We then show that this class can be easily parametrized and searched using evolutionary strategies (ES), optimizing for the final performance of the trained policy, as measured by the ground truth reward.
@@
 Analyzing the discovered algorithms, we find that the main difference between them and the baselines is that they keep optimizing the policy of the agent well after the probability of generating the chosen trajectory has surpassed the probability of generating the rejected one.
-We use this insight to design a new PO algorithm, Temporally-Aware Mirror Preference Optimization (TA-MPO), which demonstrate promising results in an LLM alignment task.
+We use this insight to design a new PO algorithm, Temporally-Aware Mirror Preference Optimization (TeMPO), which demonstrate promising results in an LLM alignment task.
 We summarize our contributions below.
@@
     We demonstrate that takeaways from our analysis on the MuJoCo setting, as well as the characteristics of the discovered PO algorithms, can be successfully transferred onto LLM tasks.
-In particular, we show that our TA-MPO algorithm significantly improves upon the baselines.
+In particular, we show that our TeMPO algorithm significantly improves upon the baselines.
 2 Preliminaries
@@
 Each pair of trajectories is ranked by a judge, who determines a chosen trajectory $\tau_{w}$ (“win”) and a rejected trajectory $\tau_{l}$ (“lose”), based on the cumulative rewards $r(\tau_{w})$ and $r(\tau_{l})$.
-Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley & Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
+Most settings assume the judge ranks trajectories according to the Bradley-Terry model (Bradley and Terry, 1952), whereby the probability of choosing $\tau_{w}$ over $\tau_{l}$ is defined as
 $\mathbb{P}(\tau_{w}\succ\tau_{l})$ $=\frac{\exp(r(\tau_{w}))}{\exp(r(\tau_{w}))+\exp(r(\tau_{l}))}=\sigma(r(\tau_{w})-r(\tau_{l})),$ (2)
@@
 When $\psi(x)=\log(x)$ and $\phi^{-1}(x)=\log(x/(1-x))$, (11) recovers the ORPO objective in (6).
-Temporally-Aware MPO (TA-MPO)
-Lastly we design a variation of MPO that gradually switches from SFT to PO. TA-MPO consists of single-phase algorithms that optimize the objective
+Temporally-Aware MPO (TeMPO)
+Lastly we design a variation of MPO that gradually switches from SFT to PO. TeMPO consists of single-phase algorithms that optimize the objective
 $\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[(1-\alpha(t))\psi(\pi(\tau_{w}))+\alpha(t)g\left(\phi^{-1}(\pi(\tau_{w}))-\phi^{-1}(\pi(\tau_{l}))\right)\big],$ (12)
@@
 $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$.
-We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient.
+We then use Adam (Kingma and Ba, 2015) to update the parameters $\zeta$ using the estimated gradient.
 In practice, to compute (19), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions.
@@
 Temporally-aware
-TA-MPO (1) 3577 $\pm$ 45 2730 $\pm$ 63 3106 $\pm$ 62 1971 $\pm$ 35
-TA-MPO (2) 3625 $\pm$ 49 3088 $\pm$ 69 3443 $\pm$ 53 1988 $\pm$ 39
-TA-MPO (3) 3352 $\pm$ 57 2256 $\pm$ 44 2725 $\pm$ 46 1923 $\pm$ 27
+TeMPO (1) 3577 $\pm$ 45 2730 $\pm$ 63 3106 $\pm$ 62 1971 $\pm$ 35
+TeMPO (2) 3625 $\pm$ 49 3088 $\pm$ 69 3443 $\pm$ 53 1988 $\pm$ 39
+TeMPO (3) 3352 $\pm$ 57 2256 $\pm$ 44 2725 $\pm$ 46 1923 $\pm$ 27
 Keep optimizing
@@
 4.3 Including temporal awareness
-We build an algorithm within the TA-MPO family using the insights gained in the previous section.
-In particular, we keep the standard SFT loss, which was rediscovered in all our experiments, and use SimPO for the PO component of TA-MPO, given its high performance and its ability to continue optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
+We build an algorithm within the TeMPO family using the insights gained in the previous section.
+In particular, we keep the standard SFT loss, which was rediscovered in all our experiments, and use SimPO for the PO component of TeMPO, given its high performance and its ability to continue optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$.
 That is, we define the objective
@@
 We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting.
-In particular, we test the TA-MPO algorithm in (13), which is designed to continue to optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, as all our discovered algorithms do.
-We evaluate TA-MPO on LLM alignment, comparing its performance against baselines for three combinations of base model and preference dataset, as shown in Table˜2.
-To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the TA-MPO objective in (13).
+In particular, we test the TeMPO algorithm in (13), which is designed to continue to optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, as all our discovered algorithms do.
+We evaluate TeMPO on LLM alignment, comparing its performance against baselines for three combinations of base model and preference dataset, as shown in Table˜2.
+To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the TeMPO objective in (13).
 We evaluate the tuned LLMs against GPT-4, using the AlpacaEval library  (Li et al., 2023) and Llama-3.1-70B-Instruct as a judge.
@@
 Refer to Appendix˜L for further details.
-Table˜2 shows the effectiveness of the TA-MPO objective in (13), which presents a high winrate for all schedules of $\alpha$.
-In particular, TA-MPO with the sigmoid schedule for $\alpha$ has the highest winrate in all tasks and the highest length controlled winrate in two out of three tasks.
-Additionally, TA-MPO requires only one stage of training, while DPO and SimPO require two, i.e.
+Table˜2 shows the effectiveness of the TeMPO objective in (13), which presents a high winrate for all schedules of $\alpha$.
+In particular, TeMPO with the sigmoid schedule for $\alpha$ has the highest winrate in all tasks and the highest length controlled winrate in two out of three tasks.
+Additionally, TeMPO requires only one stage of training, while DPO and SimPO require two, i.e.
 SFT and PO.
@@
 At the same time, the objectives discovered on TLA are only used to logits within -14 and -8, as shown in Figure˜6, and have less regular shape outside of this subset.
-In Table˜3, we report additional results that allow us to better understand the influence of different components of TA-MPO.
+In Table˜3, we report additional results that allow us to better understand the influence of different components of TeMPO.
 In particular, we have tested SimPO with $\gamma=0$, SimPO with $\gamma>0$, and SFT + SimPO with $\gamma>0$, where the last one consists of a one-step objective made of the addition between the SFT and the SimPO loss.
 Table˜3 shows that having a large $\gamma$, which encourages large updates even when $\pi(\tau_{w})>\pi(\tau_{l})$, or adding an SFT term to the PO loss are both helpful in improving performance.
-The performance of TA-MPO reported in Table˜2 is still the highest, which means that temporal awareness also contributes to a better performance.
+The performance of TeMPO reported in Table˜2 is still the highest, which means that temporal awareness also contributes to a better performance.
 We can therefore confirm that each of the insights we gained in the MuJoCo benchmark transfers to the LLM alignment setting.
@@
 Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$.
-Using this insight, we proposed a temporally-aware algorithm, TA-MPO, that avoids such loss truncation and gradually switches from the SFT step to the PO step.
+Using this insight, we proposed a temporally-aware algorithm, TeMPO, that avoids such loss truncation and gradually switches from the SFT step to the PO step.
 We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
@@
 Max length 2048
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Sections 2-5
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: In Section 5.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: See Appendix C
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Sections 4 and 5 and Appendix J.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification:
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: Appendix J
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Each quantity is reported with a standard error.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Appendix J.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification:
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49.
+Justification: This paper presents algorithms for LLM tuning.
+Misuse of these algorithms for training an LLM to produce undesirable, unethical, or harmful outputs could be possible by a user, but it is not a direct impact of our work.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification:
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: See Reference section.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification:
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification:
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification:
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification:
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.